### v5

# CLAP: Isolating Content from Style through Contrastive Learning with Augmented Prompts

Abstract
Contrastive vision-language models, such as CLIP, have garnered considerable attention for various downstream tasks, mainly due to the remarkable generalization ability of the learned features. However, the features they learn often blend content and style information, which somewhat limits their generalization capabilities under distribution shifts. To address this limitation, we adopt a causal generative perspective for multimodal data and propose contrastive learning with data augmentation to disentangle content features from the original representations. To achieve this, we begin by exploring image augmentation techniques and develop a method to seamlessly integrate them into pre-trained CLIP-like models to extract pure content features. Taking a step further, and recognizing the inherent semantic richness and logical structure of text data, we explore the use of text augmentation to isolate latent content from style features. This enables CLIP-like models’ encoders to concentrate on latent content information, refining the representations learned by pre-trained CLIP-like models. Our extensive experiments across diverse datasets demonstrate significant improvements in zero-shot and few-shot classification tasks, alongside enhanced robustness to various perturbations. These results underscore the effectiveness of our proposed methods in refining vision-language representations and advancing the state of the art in multimodal learning.
Keywords Data Augmentation  $\cdot$ Latent Variables  $\cdot$ Disentanglement
1 Introduction
Vision-language models, exemplified by CLIP [36], have garnered substantial attention due to their exceptional generalization capabilities, achieved through the learned features obtained by utilizing a cross-modal contrastive loss [36, 20, 25]. However, despite being pre-trained on extensive datasets, CLIP-like models fall short in disentangling latent content information and latent style information. Consequently, they are not immune to spurious correlations, i.e., style-related information is erroneously utilized to predict task-related labels. These limitations become evident in the presence of distribution shifts or adversarial attacks, where spurious correlations often change across different environments. For example: (1) a notable dependence on specific input text prompts has been reported for zero-shot capabilities [48, 47, 21]; (2) performance decline in few-shot scenarios has been observed [36, 13]; and (3) susceptibility to adversarial attacks has been explored [33, 45, 43].
[…]
Figure 1: Causal generative models of vision-language data. Image and text data are generated through distinct underlying deterministic processes, $\mathbf{g}_{x}$ for images and $\mathbf{g}_{t}$ for texts, derived from a unified latent space with latent content variables $\mathbf{c}$ and latent style variables $\mathbf{s}$. Latent content $\mathbf{c}$ exclusively determines the sample label $\mathbf{y}$. (a) Soft interventions on latent style variables $\mathbf{s}$ result in $\mathbf{\tilde{s}}$, subsequently generating augmented images $\mathbf{\tilde{x}}$. (b) Due to the same latent space, soft interventions on latent style variables $\mathbf{s}$ can also result in $\mathbf{\tilde{s}}$, producing augmented text $\mathbf{\tilde{t}}$. (c) A qualitative comparison of image features for zero-shot classification using "a photo of a [class]" prompts, visualized using class activation map (CAM) [32], demonstrates that while image augmentation can enhance CLIP features, the features obtained through text augmentation methods predominantly focus on the content.
Taking a causal perspective, this work begins with a simple yet effective method—image augmentation—to disentangle content and style information within the learned representations of CLIP-like models. This approach is inspired by recent advancements in theoretical developments in causal representation learning [41], which demonstrate that augmented images can be interpreted as results of soft interventions on latent style variables, as depicted in fig. 1. Such augmentation results in a natural data pair where content information remains unchanged while style information changes. Consequently, using contrastive learning, it becomes feasible to isolate the invariant content information from the variant style information. Motivated by this theoretical advancement, we propose a practical method to incorporate image augmentation into CLIP-like models to extract content information from the originally learned features. Specifically, a disentangled network is designed to fine-tune the pre-trained CLIP model using contrastive loss with image augmentation.
Despite the advancements made in disentangling content and style information from the original features learned by CLIP-like models through image augmentation, we recognize an inherent limitation: it is generally challenging to design adequate image augmentations that ensure all style factors change in an image. Theoretically, disentangling content and style information necessitates changes in all style factors [41]. However, inducing sufficient changes in latent style through image augmentation poses challenges due to the high dimensionality and complexity of style information in image data. Achieving significant style variation via artificially designed image augmentation techniques—such as transforming a photograph of a dog into a sketch while preserving content but dramatically altering style—is notably difficult.
Taking a further step, rather than relying on image augmentation, we explore the use of text augmentation to disentangle latent content and style factors. This shift is motivated by two key observations: (1) vision and language data share the same latent space, and therefore, text augmentation can also be utilized to induce changes in latent style factors; (2) text data inherently possesses high semanticity and logical structure, making it more amenable to property-wise manipulation compared to image data. Consequently, implementing sufficient style changes through text augmentation is more feasible than image augmentation, contributing to isolating content from style information (see fig. 1 for visual comparison). For instance, transforming text from "a photo of a dog" to "a sketch of a dog" is straightforward in the language modality, whereas achieving a similar transformation in image data is challenging. Inspired by these observations, we posit that introducing style variations through text augmentation, as illustrated in fig. 1, provides a more effective approach for learning vision-language content features than relying solely on image augmentation.
In summary, our contributions include:
  1. 1.
    Aimed at disentangling latent content and style factors to refine vision-language features of pre-trained CLIP-like models, we propose contrastive learning with data augmentation to fine-tune the original features from a causal perspective.
  2. 2.
    We present a novel method customized for pre-trained CLIP-like models. This method leverages a disentangled network, which is trained using contrastive learning with image augmentation, to extract latent content features from the learned features provided by the image encoder.
  3. 3.
    We propose Contrastive Learning with Augmented Prompts (CLAP) to extract latent content features from representations of CLIP-like models. It begins by training a disentangled network using the pre-trained text encoder and text augmentation. Subsequently, the trained disentangled network is transferred to the image encoder.
  4. 4.
    Experiments conducted on a large real-world dataset demonstrate the effectiveness of the proposed image and text augmentations in improving zero-shot and few-shot performance, as well as robustness against perturbations.
2 Related Work
2.1 Contrastive Vision-Language Models
Using a cross-modal contrastive loss, CLIP [36] introduced a scalable contrastive vision-language model by leveraging a large corpus of internet-sourced image-text pairs, demonstrating unprecedented zero-shot learning capabilities and exceptional generalization ability across datasets, while supporting numerous downstream tasks [38]. ALIGN [20] expanded the scale of contrastive vision-language modeling by training on up to one billion image-text pairs and integrating the vision transformer’s self-attention mechanism [11], which further enhanced performance.
Despite their successes, CLIP-like models exhibit sensitivity to input text prompts [48, 21], leading to variable performance across different prompts. Efforts to mitigate this prompt sensitivity through prompt learning and engineering [48, 21, 9, 47, 14] focus on customizing prompts for specific tasks but do not fundamentally enhance CLIP’s representations. Furthermore, CLIP-like models are vulnerable to adversarial attacks [4, 12], with current strategies [33, 45] involving adversarial-natural image pairs to improve resilience.
Our work diverges from these task-specific approaches by aiming to enhance CLIP’s representations from a disentanglement perspective, addressing the aforementioned issues inherent in CLIP-like models.
2.2 Disentangled Representation Learning
Aimed at segregating intrinsic latent factors in data into distinct, controllable representations, disentangled representation learning benefits various applications [44, 40, 24]. Specifically, in classification tasks, it has been shown that improving model performance and robustness against data distribution perturbations can be achieved by more effectively disentangling invariant content variables, without needing to fully identify all intrinsic latent variables [22, 28, 26, 27].
Within single modalities, studies such as [49] have illustrated that contrastive learning [7, 16, 18] can potentially reverse the data-generative process, aiding in the separation of representations. Furthermore, [41] suggest that image augmentation can help isolate content variables from the latent space through significant stylistic changes. [19] employs mixture techniques for data augmentation, enabling more abundant cross-modal matches.
Diverging from these methods, our approach focuses on employing text augmentation to disentangle latent content variables, introducing a unique approach to learning refined vision-language representations.
3 A Causal Generative Model for Multimodal Data
To understand pre-trained CLIP-like models, we investigate the underlying causal generative process for vision-language data. We consider the following causal generative model, as depicted in fig. 1. In the proposed model, the shared latent space governing vision and language data is divided into two distinct subspaces: one corresponding to the latent content variables $\mathbf{c}$ and the other to the latent style variables $\mathbf{s}$. The latent content variables are posited to determine the object label $\mathbf{y}$, a relationship corroborated by prior studies [22, 29, 31].
Furthermore, to elucidate the correlation between the latent style variable $\mathbf{s}$ and the object variable $\mathbf{y}$, our model incorporates the premise that the latent content variable $\mathbf{c}$ causally influences the latent style variable $\mathbf{s}$, in concordance with the principles of causal representation learning highlighted in recent literature [10, 29, 41]. Additionally, considering the diversity between image data and text data—where information in image data is typically much more detailed, while information in text data tends to be more logically structured—we posit distinct causal mechanisms for the respective generative processes. Our causal generative model is formulated using structural causal models [2]:
$\mathbf{s}:=\mathbf{g}_{s}(\mathbf{c}),~{}\mathbf{x}:=\mathbf{g}_{x}(\mathbf{c,s}),~{}\mathbf{t}:=\mathbf{g}_{t}(\mathbf{c,s}),~{}\mathbf{y}:=\mathbf{g}_{y}(\mathbf{c}).$ (1)
In Eq. 1, the style variable $\mathbf{s}$ is causally influenced by the content via $\mathbf{g}_{s}$; $\mathbf{x}$ and $\mathbf{t}$ denote visual and textual data, respectively. Both visual and textual data are causally produced by the shared latent variables $\mathbf{c}$ and $\mathbf{s}$ through distinct, reversible generative processes: $\mathbf{g}_{x}$ for images and $\mathbf{g}_{t}$ for text. The label $\mathbf{y}$ of a sample is exclusively determined by the content variable $\mathbf{c}$ via $\mathbf{g}_{y}$. For simplicity, exogenous noise variables are implicitly assumed but not explicitly represented in the formulation of the causal generative model, in line with the common understanding that each latent variable is influenced by exogenous noise.
Recent seminal work in [41] has demonstrated that the latent content variable $\mathbf{c}$ can be identified up to block identifiability (i.e., $\mathbf{c}$ can be isolated from the style variable $\mathbf{s}$) by requiring all latent style variables to change (e.g., soft interventions on all latent style variables). This change can be achieved through image augmentation; that is, the augmented image $\mathbf{\tilde{x}}$ can be interpreted as a generative result of $\mathbf{\tilde{s}}$, which is produced through soft interventions on the original latent style variables $\mathbf{s}$.
Despite this theoretical advancement, the practical implementation of this result within CLIP-like models remains unclear. In this study, we propose a practical method to disentangle content and style information within CLIP-like models by employing image augmentation, as detailed in § 4.1. Moreover, we recognize that implementing sufficient changes in all latent style variables $\mathbf{s}$ through text augmentation is more feasible than through image augmentation. Due to the high semanticity and logical structure of text data, we further explore the use of text augmentation to separate content information from style information, as discussed in § 4.2.
4 Isolating Content from Style with Data Augmentation
In this section, we propose the use of data augmentation to extract content information from the learned features in pre-trained CLIP-like models. Essentially, data augmentation facilitates the alteration of style factors while preserving content factors. Consequently, leveraging contrastive learning enables the segregation of content information from style information. We explore two distinct forms of data augmentation, namely image augmentation (§ 4.1) and text augmentation (§ 4.2).
4.1 Isolating Content from Style with Augmented Images
While recent studies (von et al., 2021) have provided theoretical assurance regarding the disentanglement of content and style through contrastive learning with data augmentation, it remains unclear how these findings can be applied to vision-language models. We extend these theoretical results to CLIP-like models as follows.
The theoretical framework suggests using InfoNCE loss [34] to extract content information, as formulated below:
$\mathcal{L}(\mathbf{f};\{\mathbf{x}_{i},\mathbf{\tilde{x}}_{i}\}_{i=1}^{b},\tau)=-\frac{1}{b}\sum\nolimits_{i=1}^{b}\log\frac{\exp\bigl{(}\langle\mathbf{f}(\mathbf{x}_{i}),\mathbf{f}(\mathbf{\tilde{x}}_{i})\rangle/\tau\bigr{)}}{\sum_{j=1}^{b}\exp\bigl{(}\langle\mathbf{f}(\mathbf{x}_{i}),\mathbf{f}(\mathbf{\tilde{x}}_{j})\rangle/\tau\bigr{)}},$ (2)
[…]
(a) Image augmentation
(b) Text augmentation
(c) Zero-shot inference
Figure 2: Refining CLIP through data augmentation. (a) Training involves a disentangled network $\mathbf{f}_{c}$, utilizing contrastive loss on original and augmented image pairs $\mathbf{x}$ and $\mathbf{\tilde{x}}$, with CLIP’s image encoder $\mathbf{f}^{*}_{x}$ having frozen gradients. (b) More efficient content feature learning is achieved through contrastive learning with augmented text prompts $\mathbf{t}$ and $\mathbf{\tilde{t}}$, using the fixed text encoder $\mathbf{f}^{*}_{t}$ of CLIP. (c) Inference stage: The trained disentangled network $\mathbf{f}^{*}_{c}$ integrates with CLIP’s text and image encoders, $\mathbf{f}^{*}_{t}$ and $\mathbf{f}^{*}_{x}$, to enable zero-shot inference for an input image $\mathbf{x}$ and class names $\mathbf{t}_{1}$ to $\mathbf{t}_{n}$.
We extend this formulation to refine pre-trained vision-language models using contrastive learning with augmented images (hereinafter referred to as "Im.Aug"). As illustrated in fig. 2(a), we train a disentangled network on top of CLIP’s pre-trained image encoder. To improve training efficiency and usability, we freeze the pre-trained image encoder. Based on the InfoNCE loss, the learning objective of Im.Aug is defined as:
$\mathbf{f}^{*}_{c}=\mathop{\mathrm{argmin}}_{\mathbf{f}_{c}}\mathop{\mathbb{E}}_{\{\mathbf{x}_{i}\}_{i=1}^{b}\in\mathcal{D}_{x}}\mathcal{L}(\mathbf{f}_{c}\circ\mathbf{f}^{*}_{x};\{\mathbf{x}_{i},\mathbf{\tilde{x}}_{i}\}_{i=1}^{b},\tau),$ (3)
where $\mathcal{D}_{x}$ denotes the training image dataset, $b$ is the batch size, and $\mathbf{f}_{c}$ is the disentangled network being trained. The pre-trained CLIP image encoder is represented by $\mathbf{f}^{*}_{x}$, with the asterisk "*" indicating that its weights are fixed. The variable $\mathbf{x}_{i}$ refers to an image sampled from $\mathcal{D}_{x}$, and $\mathbf{\tilde{x}}_{i}$ is its augmented view.
The composition of the training dataset $\mathcal{D}_{x}$, the image augmentation techniques used, the structure of the disentangled network $\mathbf{f}_{c}$, and the usage of $\mathbf{f}^{*}_{c}$ after training are detailed in the following subsections.
Data Synthesis and Image Augmentation
To generate training image data, we combine class names with various image and object attributes to create text prompts for each class. Using a stable diffusion model [39], we produce synthetic images that comprise our training dataset $\mathcal{D}_{x}$. The creation of template prompts for stable diffusion is based on attributes such as object size, color, image type, and art style. As detailed in table 1, the attributes include 10 colors and 3 object sizes, as well as 8 image types and 2 art styles. By assembling these attributes into prompts such as "a [art style] [image type] of a [object size] [object color] [class]", we generate 480 unique texts per class, from which one image per prompt is synthesized. Further details on image synthesis and examples are provided in § B.1.
Table 1: Template-based prompts. Attributes used to generate text prompts follow the structured format: "a [art style] [image type] of a [object size] [object color] [class]", where "[class]" represents the class names.
Object Color Object Size Image Type Art Style
yellow, green, black, blue, multicolored, orange, red, white, brown, purple large, small, normal sized painting, cartoon, infograph, sketch, photograph, clipart, mosaic art, sculpture realistic, impressionistic
For the image augmentation procedures, we adopt techniques commonly used in contrastive learning [7, 41, 8], specifically random cropping and color distortion.
Disentangled Network Structure
Since the training process is based on CLIP’s pre-trained lower-dimensional features, our disentangled network adopts a multi-layer perceptron (MLP) architecture. To fully benefit from the pre-trained CLIP text encoder, we construct a residual MLP featuring a zero-initialized projection, acting as the disentangled network, as depicted in fig. 3. This design enables learning directly from the pre-trained representation space, avoiding a random starting point. It is inspired by ControlNet’s zero-conv operation [46], which we adapt to a zero-linear operation within our residual MLP.
Within this architecture, the main branch includes a zero-initialized, bias-free linear layer, positioned after a SiLU activation and a normally initialized linear layer. Conventionally, the dimensions of features before the initial linear layer, between the first and second linear layers, and after the second linear layer are referred to as the input ($d_{\text{in}}$), latent ($d_{\text{mid}}$), and output ($d_{\text{out}}$) dimensions, respectively.
To rectify any mismatches between the input and output dimensions, the network employs nearest-neighbor downsampling in the shortcut path, thereby ensuring both dimensional alignment and the preservation of sharpness in the input features. During inference, a weighting parameter $\alpha>0$ is introduced to modulate the contribution of features from the main branch before integration with the input features, whereas this parameter remains fixed at 1 during training.
Inference
After training, the disentangled network $\mathbf{f}^{*}_{c}$ is utilized following CLIP’s image encoder to extract visual content features. Moreover, given that vision-language data generation is rooted in a unified latent space, as described in § 3, $\mathbf{f}^{*}_{c}$ can be seamlessly integrated with CLIP’s image and text encoders to enhance zero-shot capabilities. As shown in fig. 2(c), for an image $\mathbf{x}$, the operation is formulated as the composition function $\mathbf{f}^{*}_{c}\circ\mathbf{f}^{*}_{x}(\mathbf{x})$, and similarly, for a text $\mathbf{t}$, as $\mathbf{f}^{*}_{c}\circ\mathbf{f}^{*}_{t}(\mathbf{t})$. This integration preserves CLIP’s zero-shot functionality while providing refined features through improved disentanglement of content.
Figure 3: Structure of the disentangled network. The architecture encompasses a residual block featuring a zero-initialized, bias-free linear layer to initiate optimization from the input feature space. When the input and output dimensions differ, a downsampling operation is employed to ensure alignment. During inference, a scalar parameter $\alpha$ balances the contributions of the main branch and the input features prior to their combination.
4.2 Isolating Content from Style with Augmented Prompts
Despite progress in disentangling content and style via image augmentation, adequately altering all style factors in an image remains challenging due to the high dimensionality and complexity of style information in images. Achieving substantial style changes through augmentation—essential for complete disentanglement [41]—is difficult with existing image augmentation techniques. On the contrary, text data inherently possesses high semanticity and logical structure, making it more amenable to property-wise manipulation compared to image data. To further explore the disentanglement of content, we propose Contrastive Learning with Augmented Prompts (CLAP).
As depicted in fig. 2(b), CLAP employs an InfoNCE loss to train a disentangled network atop CLIP’s pre-trained text encoder, keeping the encoder’s gradients fixed—similar to Im.Aug. Leveraging the simpler structure of text, the template-based prompts previously utilized for synthesizing images now serve as the training text dataset, denoted by $\mathcal{D}_{t}$. Utilizing the same disentangled network as in Im.Aug, the learning objective of CLAP is defined as:
$\mathbf{f}^{*}_{c}=\mathop{\mathrm{argmin}}_{\mathbf{f}_{c}}\mathop{\mathbb{E}}_{\{\mathbf{t}_{i}\}_{i=1}^{b}\in\mathcal{D}_{t}}\mathcal{L}(\mathbf{f}^{*}_{c}\circ\mathbf{f}^{*}_{t}(\mathbf{t});\{\mathbf{t}_{i},\mathbf{\tilde{t}}_{i}\}_{i=1}^{b},\tau)+\lambda\mathcal{L}(\mathbf{f}^{*}_{c}\circ\mathbf{f}^{*}_{t}(\mathbf{t});\{\mathbf{t}^{c}_{i},\mathbf{\tilde{t}}_{i}\}_{i=1}^{b},1),$ (4)
where $\mathbf{f}^{*}_{t}$ denotes the pre-trained CLIP text encoder. The term $\mathbf{t}_{i}$ refers to a text prompt from $\mathcal{D}_{t}$, and $\mathbf{\tilde{t}}_{i}$ represents its augmented view, produced via prompt augmentation techniques. On the right-hand side of the equation, $\mathbf{t}_{i}^{c}$ denotes the class name associated with the text prompt $\mathbf{t}_{i}$. This strategy aims to enhance variation between prompt pairs, especially in cases where the text dataset $\mathcal{D}_{t}$ contains a limited number of samples. The hyperparameter $\lambda$ controls the relative weight of the second term in the overall loss function. All other symbols in Eq. 4 retain their definitions as described earlier.
After training, the learned disentangled network is seamlessly integrated with both of CLIP’s encoders to extract content representations, as shown in fig. 2(c).
Prompt Augmentation
To ensure text prompts undergo stylistic changes without compromising their content, we develop specific augmentation techniques for synthetic text prompts. Drawing inspiration from Easy Data Augmentation (EDA) [42], we adapt the Random Deletion (RD) and Random Swap (RS) techniques, customizing them for our structured prompt format. To avoid inadvertently altering the content—such as by introducing new object names or modifying the core semantics—we exclude random insertions and replacements.
Our primary augmentation techniques include Object Size Deletion (OSD), Object Color Deletion (OCD), Image Type Deletion (ITD), Art Style Deletion (ASD), and Swapping Prompt Order (SPO). Each is applied with a fixed probability, as detailed in table 2. Additionally, for downstream datasets with few categories, we enrich the training set population using an additional augmentation method named IGN (Inserting Gaussian Noise). Following initialization protocols from prompt learning [48, 47], we insert zero-mean Gaussian noise with a standard deviation of 0.02 and noise length of 4 into the tokenized prompts.
Intuitively, these prompt augmentation methods parallel random masking techniques used in image augmentation [17, 6]. However, prompt augmentations are often more effective and precise than their image-based counterparts. This is because prompt augmentations can directly target and remove specific style elements without affecting core content, whereas image masking operates at the pixel or patch level and may inadvertently distort content or fail to sufficiently alter style.
Table 2: Prompt augmentation techniques. Various augmented views are generated from an original text prompt using specific techniques: OSD (Object Size Deletion), OCD (Object Color Deletion), ITD (Image Type Deletion), ASD (Art Style Deletion), and SPO (Swapping Prompt Order).
Prompt Variant OSD OCD ITD ASD SPO
a realistic painting of a large red car a realistic painting of a red car a realistic painting of a large car a realistic of a large red car a painting of a large red car a large red car in a realistic painting
5 Experiments
We conduct three primary experiments to assess our method:
  1. 1.
    Zero-shot evaluation with diverse prompts to gauge performance and robustness under prompt perturbations;
  2. 2.
    Linear probe tests on few-shot samples to evaluate the efficacy of the learned representations in few-shot settings;
  3. 3.
    Adversarial attack assessments on zero-shot and one-shot classifiers to determine their resistance to adversarial threats.
We further conduct an ablation study on hyperparameters, explore the impact of different prompt augmentation combinations and various sources of training prompts on CLAP’s performance, and replicate experiments across different CLIP model sizes.
5.1 Experimental Setup
Implementation
Im.Aug and CLAP are implemented using the ViT-B/16 CLIP model and executed on an NVIDIA RTX 3090 GPU. To ensure reproducibility, the random seed for all stochastic processes is fixed at 2024. Additional implementation details are provided in § A.1.
Datasets
CLAP is evaluated on four multi-domain datasets to assess its performance across diverse environments: PACS [23] (4 domains, 7 categories), VLCS [1] (4 domains, 5 categories), OfficeHome [37] (4 domains, 65 categories), and DomainNet [35] (6 domains, 345 categories). For conciseness, we present average results across domains for each dataset. Detailed results for individual domains are provided in § A.4.
Compute Efficiency
CLAP demonstrates faster convergence and shorter training times compared to Im.Aug. Specifically, CLAP completes training on the PACS and VLCS datasets in approximately 11 minutes, OfficeHome in about 14 minutes, and DomainNet in roughly 47 minutes. In contrast, Im.Aug requires around 16 minutes for PACS and VLCS, 50 minutes for OfficeHome, and 3.3 hours for DomainNet. Both Im.Aug and CLAP retain CLIP’s inference efficiency due to the disentangled network’s compact two-layer MLP structure.
5.2 Main Results
Zero-Shot Performance
To assess zero-shot capabilities, CLAP is evaluated using three specific fixed prompts: ZS(C), which uses only the class name as "[class]"; ZS(PC), formatted as "a photo of a [class]"; and ZS(CP), structured as "a [class] in a photo". To further examine robustness, we introduce a dynamic prompt, ZS(NC), structured as "[noise][class]", where "[noise]" refers to Gaussian noise with mean 0 and standard deviation 0.02.
As shown in table 3, CLAP outperforms both CLIP and Im.Aug across all prompts and datasets. Unlike the uniform improvement achieved by CLAP, Im.Aug yields inconsistent results. Notably, CLAP exhibits a substantial advantage over CLIP when evaluated with the dynamic ZS(NC) prompt, demonstrating its effectiveness in improving zero-shot performance through enhanced content disentanglement.
Table 3: Zero-shot performance across prompt variants. Fixed prompts: "C" for [class], "CP" for a [class] in a photo, "PC" for a photo of a [class]; dynamic prompt "NC" is [noise][class] with Gaussian noise. CLAP consistently outperforms CLIP, while image augmentation (Im.Aug) shows mixed results.
Prompt Method        PACS      VLCS OfficeHome DomainNet Overall
ZS(C) CLIP 95.7 76.4 79.8 57.8 77.4
Im.Aug 96.5 79.5 77.0 51.5 76.1
CLAP 97.2 82.6 81.0 58.7 79.9
ZS(CP) CLIP 95.2 82.0 79.5 57.0 78.4
Im.Aug 96.3 82.9 75.8 50.7 76.4
CLAP 97.3 83.4 80.5 58.0 79.8
ZS(PC) CLIP 96.1 82.4 82.5 57.7 79.7
Im.Aug 96.5 83.0 78.6 51.6 77.4
CLAP 97.2 83.4 83.0 59.0 80.6
ZS(NC) CLIP 90.8 68.3 71.5 51.0 70.4
Im.Aug 94.8 73.1 67.5 44.0 69.9
CLAP 97.2 81.0 73.5 52.6 76.1
To evaluate robustness to prompt perturbations, we analyze the variance in zero-shot performance across prompts using the range ($R$) and standard deviation ($\delta$) from ZS(C), ZS(CP), and ZS(PC). We also measure the performance drop ($\Delta_{(NC)}$) from ZS(C) to ZS(NC) to assess resilience to noisy prompts.
As shown in table 4, CLAP significantly reduces zero-shot variance, exhibiting lower $\delta$ and $R$ values, and a smaller drop in performance under noise, compared to both Im.Aug and CLIP. While Im.Aug offers modest variance reduction, it is clearly less effective than CLAP. These results demonstrate CLAP’s superior robustness in maintaining consistent performance across prompt variations.
Table 4: Prompt robustness analysis. CLAP more effectively reduces zero-shot performance variance across prompts than image augmentation. $R$ and $\delta$ denote the range and standard deviation across ZS(C), ZS(CP), and ZS(PC); $\Delta_{(NC)}$ indicates the performance drop from ZS(C) to ZS(NC). Lower values imply greater robustness.
Metric Method        PACS      VLCS OfficeHome DomainNet Overall
$R$ CLIP 0.9 6.1 3.1 0.8 2.7
Im.Aug 0.1 3.6 2.8 0.9 1.9
CLAP 0.1 0.8 2.5 1.0 1.1
$\delta$ CLIP 0.4 2.8 1.4 0.4 1.2
Im.Aug 0.1 1.7 1.2 0.4 0.8
CLAP 0.0 0.4 1.1 0.4 0.5
$\Delta_{(NC)}$ CLIP 4.9 8.1 8.3 6.8 7.0
Im.Aug 1.6 6.4 9.5 7.5 6.3
CLAP 0.0 1.6 7.5 6.1 3.8
Few-Shot Performance
We conduct evaluations using 1-shot, 4-shot, 8-shot, and 16-shot linear probes across each domain within the four datasets. As illustrated in fig. 4, CLAP significantly outperforms both CLIP and Im.Aug in few-shot learning scenarios. Notably, in the 1-shot setting, CLAP achieves performance improvements over the linear-probe CLIP model by margins of +10%, +3.5%, +2.5%, and +1.5% on the PACS, VLCS, OfficeHome, and DomainNet datasets, respectively. These improvements are especially pronounced when compared to the gains observed with Im.Aug, highlighting CLAP’s superior effectiveness in few-shot learning. Detailed quantitative results are provided in § A.4.
Figure 4: Few-shot linear probe comparisons of image-encoder features show that CLAP enhances CLIP’s few-shot performance more effectively than Im.Aug. "ZS" indicates zero-shot performance using a [class] prompt.
Adversarial Performance
To evaluate adversarial robustness, we test both zero-shot (ZS(C)) and one-shot classifiers against three popular adversarial attack methods: FGSM [15], PGD [30], and CW [5]. Adversarial samples are generated with 1 iteration for FGSM and 20 iterations for both PGD and CW, using an $\epsilon$ of 0.031.
As shown in table 5, classifiers leveraging CLAP representations exhibit significantly greater robustness than those based on CLIP. Across the four datasets, CLAP’s zero-shot and one-shot classifiers outperform CLIP by +7.6% and +8.5% against FGSM, +1.0% and +11.7% against PGD-20, and +1.1% and +2.3% against CW-20, respectively. These gains surpass the improvements achieved by Im.Aug: +4.4% and +4.6% (FGSM), +0.3% and +6.2% (PGD-20), and 0% and +1.3% (CW-20). These results indicate that CLAP enhances robustness against adversarial attacks in both zero-shot and one-shot scenarios.
Table 5: Adversarial robustness evaluation. Both Im.Aug and CLAP improve the robustness of CLIP’s zero-shot ([class]) and 1-shot classifiers under adversarial attacks (FGSM, PGD-20, CW-20). CLAP consistently achieves the highest resilience across all datasets.
Setting Method FGSM PGD-20 CW-20 Avg.
PACS VLCS O.H. D.N. PACS VLCS O.H. D.N. PACS VLCS O.H. D.N.
ZS(C) CLIP 86.8 65.6 57.9 22.5 29.1 2.0 10.1 10.7 27.4 1.5 7.4 7.6 29.2
Im.Aug 88.0 69.6 55.1 37.9 31.3 2.1 10.4 9.0 29.4 1.7 7.0 5.8 31.1
CLAP 88.7 71.9 58.5 44.2 30.8 3.2 10.6 11.2 29.8 2.3 8.1 8.0 32.7
1-shot CLIP 66.7 45.2 34.3 22.5 34.8 16.0 5.6 11.3 18.9 0.7 4.5 3.2 23.7
Im.Aug 79.4 47.1 37.1 23.5 55.2 16.1 8.5 12.5 23.2 0.9 5.1 3.4 28.0
CLAP 89.6 52.2 37.1 23.9 73.4 21.2 7.4 12.5 27.0 1.1 5.0 3.5 31.9
5.3 More Analysis
t-SNE Visualization
In our t-SNE visualizations, we examine the representations of CLIP, Im.Aug, and CLAP for all images within the Art Painting domain of the PACS dataset. fig. 5 shows that CLAP’s image representations display a marked inter-class separation and tighter intra-class clustering than those of CLIP and Im.Aug. This observation suggests that CLAP’s representations are more closely tied to content information and less influenced by style information, in contrast to the other two.
(a) CLIP
(b) Im.Aug
(c) CLAP
Figure 5: t-SNE visualizations of all images in the Art Painting of PACS dataset show CLAP outperforms the original CLIP and Im.Aug, with clearer inter-class distinctions and tighter intra-class clusters.
Ablations
In fig. 6, we assess the zero-shot capabilities of our model using two distinct prompts, ZS(C) and ZS(PC), on the VLCS dataset. This analysis forms part of an ablative study aimed at understanding the influence of various hyper-parameters on model performance. Specifically, we examine: the dimensions of the latent layer within the MLP of the disentangled network, as illustrated in fig. 6(a); the temperature parameter ($\tau$) in the loss function, as depicted in fig. 6(b); and the weight coefficient ($\alpha$) during the inference stage, as shown in fig. 6(c). Our findings indicate that CLAP consistently enhances zero-shot performance across all tested configurations for both prompts, while also significantly reducing the gap between the performances elicited by each prompt. These results underscore the efficacy of CLAP in accommodating a wide range of hyper-parameters.
(a) Latent dimensions
(b) Temperature $\tau$ in loss
(c) Inference weight $\mathrm{log}_{10}(\alpha)$
Figure 6: We conduct ablative study on hyper-parameter choices on the VLCS dataset, including latent dimensions, $\tau$ values, and $\alpha$ values during the inference stage. CLAP continuously enhance CLIP’s performance throughout the tested values.
Prompt Augmentation Combinations
We explore diverse combinations of our tailored prompt augmentation techniques and examine the impact of Easy Data Augmentation (EDA) methods [42] on the VLCS dataset. All evaluated techniques demonstrate that CLAP provides consistent improvements over CLIP. Further details are presented in § A.2.
Prompt Sources
We assess the influence of different training prompt sources on performance, using the VLCS dataset and incorporating EDA techniques. The evaluated prompt formats include: (1) our template-based prompts; (2) LLM-generated prompts from ChatGPT-3.5 [3], with generation details in § B.2; (3) prompts of the form "a [random] style of [class]," where "[random]" is sampled from a random word generator [^1] https://github.com/vaibhavsingh97/random-word; and (4) prompts generated by the PromptStyler method [9]. The results suggest that training with simpler, more structured prompts tends to yield better performance. Detailed quantitative results are reported in § A.3.
Experiments on Different Model Scales
We repeat our zero-shot evaluations using larger pre-trained CLIP models, specifically ViT-L/14 and ResNet50x16. Across both model scales, CLAP consistently improves zero-shot performance while also reducing variance. These findings reaffirm CLAP’s scalability and robustness in enhancing CLIP representations. Quantitative results are provided in §§ B.3 and B.4.
6 Conclusion
To enhance the representations of pre-trained CLIP-like models, this study investigates the disentanglement of latent content variables. Adopting a causal perspective on the generative processes underlying vision-language data, we demonstrate that training a disentangled network in one modality enables effective disentanglement across both modalities. Leveraging the semantic richness and logical structure of text data, we find that disentanglement is more readily achieved in the language modality through targeted text augmentation.
Based on these insights, we propose CLAP (Contrastive Learning with Augmented Prompts), a method designed to acquire disentangled vision-language content features. Extensive experiments validate the effectiveness of CLAP, showing substantial improvements in zero-shot and few-shot classification tasks, along with enhanced robustness to input perturbations. We hope this work will motivate further research into disentangled representation learning within multimodal frameworks.
Appendix A More on Experiments with ViT-B/16
A.1 Implementation Details
This section provides detailed implementation settings for our experiments using the CLIP pre-trained ViT-B/16 model.
Network.
The network’s output dimension is set to match the 512-dimensional CLIP feature space, thereby eliminating the need for feature downsampling. The latent dimension is tailored for each dataset: 256 for PACS, 448 for OfficeHome, and 512 for VLCS and DomainNet, to accommodate differences in category counts and dataset complexity. The weight parameter $\alpha$ is adjusted per dataset as follows: 0.208 for PACS, 0.056 for VLCS, 0.14 for OfficeHome, and 0.2 for DomainNet. During training, $\alpha$ is consistently set to 1.
Training CLAP.
We use consistent training parameters across all datasets. The Adam optimizer is employed with a learning rate of 0.0001. Training is limited to 8,000 steps, with the average loss evaluated every 480 steps, and early stopping applied if the loss does not decrease by at least 0.01 over five checkpoints. Batch sizes are set to 8 for PACS and VLCS, 96 for OfficeHome, and 384 for DomainNet. The temperature parameter $\tau$ is configured to 0.5 for PACS and VLCS, and 0.3 for OfficeHome and DomainNet. The loss coefficient $\lambda$ is set to 1 for PACS and VLCS, and 0.0001 for OfficeHome and DomainNet, reflecting the fewer number of classes in the former two datasets.
Prompt augmentations OSD, OCD, and SPO are applied across all datasets with a probability of 0.5. For PACS and VLCS, Gaussian noise (mean 0, standard deviation 0.02) is randomly inserted at the beginning, middle, or end of the augmented prompts to diversify the training samples. For linear probe evaluations in the few-shot setting, we use L2 normalization and cross-entropy loss, training over 1,000 epochs with a batch size of 32. Early stopping is used with a patience of 10 epochs and a loss improvement threshold of 0.001.
Training Im.Aug.
We train a disentangled network using image augmentations and the InfoNCE loss with a temperature parameter $\tau$ set to 0.5. Image augmentations include random cropping (scale $\in[0.64,1.0]$) and color distortion (brightness = 0.5, hue = 0.3), each applied with a probability of 0.5. All other training and inference configurations for Im.Aug are consistent with those used for CLAP across the datasets.
A.2 Prompt Augmentation Combinations
In table 6, we evaluate various combinations of our tailored prompt augmentation techniques alongside Easy Data Augmentation (EDA) [42] on the VLCS dataset. Across all configurations, CLAP consistently enhances CLIP’s performance while effectively reducing performance variance across different prompt conditions.
Among the tested combinations, OSD+OCD+SPO+IGN (①②⑤⑥) yields the highest average zero-shot accuracy and the lowest variance, outperforming all other augmentation strategies, including EDA. Notably, even without the inclusion of random noise (IGN), CLAP achieves substantial improvements over CLIP, particularly in mitigating performance degradation under noisy prompt conditions, as reflected by a significantly lower $\Delta_{(NC)}$.
Table 6: Effect of prompt augmentation combinations on VLCS. We assess the impact of various augmentations: OSD (①), OCD (②), ITD (③), ASD (④), SPO (⑤), and IGN (⑥). ZS(Avg.) denotes the average zero-shot accuracy across four different inference prompts. CLAP improves both accuracy and robustness, with the best results achieved using the combination ①②⑤⑥.
Metric CLIP (base) EDA ①②③④⑤⑥ ①②③④⑤ ①②③④⑥ ①②③④ ③④⑤ ①②⑤⑥
ZS(Avg.) (↑) 77.3 81.6 82.0 80.1 82.0 79.6 82.1 82.6
$R$ (↓) 6.1 1.9 1.2 2.5 0.9 3.2 1.6 0.8
$\delta$ (↓) 2.8 0.9 0.6 1.2 0.4 1.5 0.7 0.4
$\Delta_{(NC)}$ (↓) 8.1 2.3 1.7 3.0 1.8 3.4 2.0 1.6
A.3 Prompt Sources
In table 7, we evaluate the impact of different training prompt formats, sourced from various synthetic origins, on the performance of CLAP using the VLCS dataset. All experiments in this subsection are conducted with EDA (Easy Data Augmentation) techniques applied during training.
The evaluated prompt formats include: Template, representing our structured template-based prompts; LLM, referring to prompts generated using ChatGPT-3.5 [3] (see § B.2 for generation details); Random, which uses the format "a [random] style of [class]", where "[random]" is filled using a random word generator; and Prm.Stl., which denotes vectorized prompts produced by the PromptStyler method [9].
Table 7: Effect of training prompt sources on CLAP performance (VLCS). We train CLAP using EDA-augmented prompts from various sources. All prompt types improve CLIP’s baseline zero-shot performance. Simpler prompt formats—Random and Template—achieve the highest accuracy and robustness.
Metric CLIP (base) LLM Random Prm.Stl. Template
ZS(Avg.) (↑) 77.3 78.2 81.6 81.2 81.6
$R$ (↓) 6.1 3.2 0.7 2.7 1.9
$\delta$ (↓) 2.8 1.5 0.3 1.2 0.9
$\Delta_{(NC)}$ (↓) 8.1 3.3 2.3 3.0 2.3
The results indicate that CLAP trained with any of the evaluated prompt sources leads to consistent improvements over the original CLIP model. Interestingly, despite the sophistication of LLM-generated and PromptStyler-based prompts, the simpler Random and Template prompts outperform them in both accuracy and robustness.
It is worth noting, however, that while these alternative prompt formats paired with EDA enhance zero-shot performance, they do not surpass the performance achieved by prompt augmentations specifically optimized for the template-based format (as reported in § A.2).
A.4 Detailed Results on ViT-B/16
Details on Zero-Shot Evaluations
We present domain-level zero-shot performance using various prompts across each dataset in table 8. CLAP consistently improves upon CLIP across all prompt variants and domains. Notably, as CLAP exclusively uses text data for training, it retains CLIP’s domain generalization capabilities—an outcome of CLIP’s large-scale training corpus. Rather than compromising this generalization, CLAP enhances it through more effective disentanglement of content information, leading to consistent zero-shot gains across all evaluated domains.
Table 8: Domain-level zero-shot results with ViT-B/16. We report top-1 accuracy (%) using four different prompts: ZS(C), ZS(CP), ZS(PC), and ZS(NC), across dataset domains. CLAP improves zero-shot performance in all settings.
Datasets Domain-level avg. top-1 acc. (%) of zero-shot performance usig ViT-B/16 (↑)
ZS(C) ZS(CP) ZS(PC) ZS(NC)
CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP
PACS A. 96.4 96.9 97.5 93.4 97.0 97.6 97.4 97.6 97.6 87.8 93.5 97.1
C. 98.9 99.0 98.9 99.0 99.2 99.0 99.1 99.0 98.9 95.4 97.6 98.8
P. 99.9 99.9 99.9 99.3 99.6 99.9 99.9 99.9 99.9 93.1 99.0 99.9
S. 87.7 90.1 92.5 89.2 89.6 92.5 88.1 89.4 92.3 87.1 89.3 93.1
VLCS C. 99.7 99.8 99.9 99.9 99.9 99.9 99.9 99.9 99.9 87.0 96.0 99.9
L. 61.8 66.2 67.7 69.9 70.4 70.4 70.2 70.2 70.7 55.9 59.9 65.9
S. 70.1 74.8 78.0 73.3 76.0 77.2 73.6 76.4 76.9 61.4 66.2 75.3
V. 73.9 77.1 84.9 84.8 85.4 86.0 86.1 85.6 86.2 68.9 70.3 82.9
O.H. A. 80.5 79.0 81.8 80.1 76.0 81.6 83.2 78.7 83.2 73.0 69.2 73.6
C. 64.6 59.6 66.4 63.7 58.9 65.4 68.1 61.9 69.0 57.0 52.0 60.4
P. 86.3 83.6 87.5 86.6 83.4 87.2 89.1 86.6 89.7 77.2 72.3 78.9
R. 88.0 85.9 88.5 87.6 84.8 87.7 89.8 87.2 90.0 79.0 76.5 81.1
D.N. C. 71.0 64.3 71.9 70.5 62.1 72.0 71.3 63.4 72.8 63.2 53.9 64.6
I. 48.6 40.5 50.6 47.7 40.7 49.5 47.8 40.0 50.5 42.9 35.0 45.1
P. 66.6 59.1 67.7 66.0 59.0 67.3 66.5 59.8 68.4 57.2 50.4 59.4
Q. 14.9 12.4 15.2 13.3 11.5 13.8 14.1 11.8 14.3 12.0 9.2 13.1
R. 82.6 76.6 83.1 82.2 75.8 82.2 83.4 78.2 83.7 75.2 67.9 75.6
S. 63.1 56.1 63.7 62.2 55.0 63.1 63.4 56.4 64.4 55.7 47.5 57.6
Table 9: Domain-level few-shot results with ViT-B/16. We report top-1 accuracy (%) for 1-shot through 32-shot settings across dataset domains. CLAP outperforms both CLIP and Im.Aug, particularly in low-shot scenarios.
Datasets Domain-level avg. top-1 acc. (%) of few-shot performance of ViT-B/16 (↑)
1-shot 4-shot 8-shot 16-shot 32-shot
CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP
PACS A. 79.5 84.1 94.5 92.4 96.4 97.2 95.1 97.2 98.4 97.9 98.1 98.4 98.8 99.1 98.9
C. 86.7 96.1 98.3 96.8 98.6 99.2 98.8 98.9 99.3 99.5 99.2 99.5 99.6 99.6 99.6
P. 97.4 99.8 99.9 99.6 99.8 99.9 99.9 99.9 99.9 99.8 99.9 99.9 99.9 99.9 99.9
S. 75.1 80.0 87.3 91.1 92.3 92.5 92.3 92.3 92.9 92.4 92.6 93.1 93.9 94.2 94.1
VLCS C. 99.2 99.7 99.8 99.9 99.8 99.9 99.8 99.7 99.9 99.7 99.9 99.9 99.9 100.0 99.9
L. 41.3 41.3 41.1 56.7 57.0 59.8 46.2 36.8 48.3 59.4 60.4 62.6 60.4 60.7 61.9
S. 45.3 46.1 50.8 61.9 63.7 69.0 67.4 67.7 71.3 75.9 76.8 80.9 77.4 78.6 81.0
V. 50.9 53.4 59.0 64.5 66.7 76.1 75.4 74.1 78.7 72.6 73.9 77.7 85.7 86.1 87.9
O.H. A. 42.6 45.1 43.9 76.8 77.6 77.7 84.8 86.0 85.5 91.8 92.1 92.1 97.4 97.5 97.5
C. 40.1 45.0 43.8 69.9 70.2 70.5 75.8 75.9 76.6 81.6 81.6 81.6 89.0 89.0 89.2
P. 70.2 73.3 73.4 89.7 90.3 90.2 93.8 93.7 93.9 95.7 95.7 95.8 97.7 97.6 97.6
R. 58.4 59.3 59.4 81.7 83.1 82.9 89.7 89.5 89.9 92.9 92.7 93.2 95.8 95.8 95.8
D.N. C. 42.1 43.6 43.8 66.8 67.5 67.8 74.2 74.3 74.6 78.5 78.6 78.8 82.8 82.8 82.7
I. 19.5 20.8 21.0 38.5 39.3 39.7 46.7 47.0 47.3 53.2 53.2 53.6 60.0 59.9 60.1
P. 32.1 33.5 34.2 60.5 60.9 61.5 68.0 68.0 68.7 72.5 72.6 73.0 76.7 76.6 76.8
Q. 15.2 15.3 15.3 30.0 29.6 29.9 37.1 36.4 36.8 43.8 43.4 43.5 49.4 49.1 49.0
R. 50.8 52.1 52.7 76.7 77.0 77.6 81.7 81.9 82.2 84.0 83.9 84.3 85.9 85.9 86.0
S. 33.1 33.9 34.8 56.2 56.6 57.2 62.9 62.9 63.7 67.8 67.7 68.1 72.5 72.3 72.6
Details on Few-Shot Evaluations
We present detailed quantitative results of few-shot performance in table 9. CLAP consistently enhances few-shot capabilities, demonstrating improvements across all evaluated domains and test datasets.
Details on Adversarial Evaluations
In table 10, we report domain-level adversarial performance for PACS, VLCS, OfficeHome, and DomainNet. CLAP improves both zero-shot and one-shot robustness across all domains. Although Im.Aug strengthens one-shot robustness, its impact on zero-shot adversarial performance remains inconsistent.
Details on Ablative Analysis
table 11 presents the detailed results of zero-shot performance using various combinations of prompt augmentation strategies. Additionally, table 12 summarizes the effects of key hyperparameters—including temperature $\tau$, latent dimension size, and weight coefficient $\alpha$—evaluated across individual domains. These findings confirm that CLAP maintains effectiveness across a broad range of hyperparameter configurations.
Table 10: Domain-level adversarial results (ViT-B/16). Average top-1 accuracy (%) under adversarial attacks: FGSM, PGD-20, and CW-20, evaluated in both zero-shot (ZS-C) and one-shot settings across domains of PACS, VLCS, OfficeHome, and DomainNet.
Datasets FGSM PGD-20 CW-20
ZS-C 1-shot ZS-C 1-shot ZS-C 1-shot
CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP
PACS A 76.3 79.3 79.3 61.2 78.0 87.3 1.7 2.2 1.8 16.0 42.1 63.1 1.5 2.0 2.3 0.5 1.1 1.7
C 94.9 95.0 94.0 66.5 84.2 95.1 33.3 37.7 35.6 33.3 57.2 86.1 28.8 34.0 33.2 11.9 23.6 31.8
P 91.6 90.3 91.7 67.4 80.8 92.1 5.7 7.0 6.7 27.1 55.0 69.8 4.7 4.9 5.8 0.7 2.7 4.1
S 84.5 87.5 89.8 71.6 74.6 83.8 75.8 78.4 79.2 63.0 66.3 74.6 74.5 76.8 77.9 62.7 65.4 70.3
VLCS C 55.3 53.8 55.5 25.8 28.8 25.3 4.4 5.1 4.7 2.0 5.2 2.5 2.9 3.1 3.5 0.7 1.2 1.0
L 49.4 45.5 50.6 27.0 32.6 30.4 15.2 14.9 16.0 6.4 8.9 8.0 12.4 11.2 13.0 6.1 8.3 7.7
S 61.7 58.1 62.5 48.0 46.9 51.6 13.2 13.9 14.0 8.6 10.7 10.0 9.2 8.8 10.2 8.3 7.9 8.4
V 65.3 63.2 65.6 36.5 40.1 41.0 7.5 7.9 7.9 5.3 9.4 8.9 5.2 4.8 5.6 2.9 2.8 2.9
O.H. A 55.3 53.8 55.5 25.8 28.8 25.3 4.4 5.1 4.7 2.0 5.2 2.5 2.9 3.1 3.5 0.7 1.2 1.0
C 49.4 45.5 50.6 27.0 32.6 30.4 15.2 14.9 16.0 6.4 8.9 8.0 12.4 11.2 13.0 6.1 8.3 7.7
P 61.7 58.1 62.5 48.0 46.9 51.6 13.2 13.9 14.0 8.6 10.7 10.0 9.2 8.8 10.2 8.3 7.9 8.4
R 65.3 63.2 65.6 36.5 40.1 41.0 7.5 7.9 7.9 5.3 9.4 8.9 5.2 4.8 5.6 2.9 2.8 2.9
D.N. C 57.8 50.9 58.8 33.3 34.3 35.0 21.6 18.7 22.8 18.4 19.6 20.0 15.8 12.5 16.6 7.0 7.5 7.8
I 35.8 28.0 37.0 12.2 13.3 13.2 6.1 3.7 6.7 4.6 5.3 5.1 3.3 1.9 3.7 0.9 0.9 0.9
P 43.9 39.0 44.3 18.4 20.6 20.3 3.1 2.8 3.3 8.6 10.4 9.9 1.8 1.3 1.9 0.3 0.3 0.3
Q 12.9 10.3 13.2 10.9 10.8 11.1 8.4 6.8 8.6 5.4 5.4 5.6 7.1 5.4 7.4 4.9 4.8 5.1
R 62.1 55.9 62.4 34.5 35.9 36.5 7.1 6.5 7.5 17.6 19.7 19.6 4.5 3.4 4.7 1.2 1.4 1.4
S 49.1 43.3 49.7 25.7 26.0 27.5 17.8 15.5 18.6 13.6 14.4 15.1 13.4 10.2 13.9 5.0 5.2 5.6
Appendix B Data Synthesis
B.1 Synthetic Image Generation
We utilize the Stable Diffusion v2.1 model [39] for generating synthetic image data used in our comparative experiments. Specifically, we adopt the implementation provided via the Stable Diffusion v2-1 Model Card on Hugging Face [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1. For each class across the four benchmark datasets, a total of 480 images are synthesized by applying our structured template prompts as input to the diffusion model. All generated images are rendered at a resolution of $512\times 512$ pixels. Representative examples of the synthetic images, along with their corresponding text prompts, are presented in fig. 7.
Figure 7: Representative examples of synthetic images generated using Stable Diffusion v2.1, paired with their associated template-based prompts.
Table 11: Zero-shot performance on VLCS across varied augmentation combinations and prompt sources. We evaluate average top-1 accuracy (%) for each domain using CLIP and CLAP variants trained with different prompt augmentations: ① Object Size Deletion, ② Object Color Deletion, ③ Image Type Deletion, ④ Art Style Deletion, ⑤ Prompt Order Swapping, and ⑥ Gaussian Noise Addition. Rightmost columns show EDA-based augmentations: LLM, Random, PromptStyler (Pr.St.), and Template.
Method Domain Avg. top-1 accuracy (%) (↑) on VLCS
CLIP (base) ①②③ ①②③ ①②③ ①②③ ③④⑤ ①②⑤ LLM Random Pr.St. Template
④⑤⑥ ④⑤ ④⑥
ZS(C) C 99.7 99.9 99.8 99.9 99.8 99.9 99.9 97.9 99.7 99.9 99.9
L 61.8 66.6 62.3 67.0 62.2 66.2 67.7 66.2 69.0 67.3 66.5
S 70.1 78.1 75.5 78.0 74.3 78.5 78.0 73.2 76.9 73.5 76.9
V 73.9 82.8 80.6 83.2 79.3 82.7 84.9 72.6 81.8 81.8 81.9
ZS(CP) C 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.8 99.9 99.9 99.9
L 69.9 69.3 67.9 69.6 68.4 70.0 70.4 69.3 70.4 71.2 69.7
S 73.3 77.6 76.4 76.7 75.9 78.8 77.2 76.2 75.2 75.1 78.0
V 84.8 85.3 84.0 85.3 84.2 85.1 86.0 77.0 84.2 86.0 84.6
ZS(PC) C 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9
L 70.2 70.0 68.0 70.1 68.5 70.0 70.7 67.5 70.6 71.8 70.0
S 73.6 76.6 75.6 76.0 74.8 77.8 76.9 76.9 75.1 74.9 78.2
V 86.1 85.7 84.7 85.7 84.5 85.5 86.2 78.2 84.6 86.8 84.8
ZS(NC) C 87.0 99.8 99.6 99.8 99.4 99.7 99.9 95.3 98.6 99.6 99.8
L 55.9 65.2 61.3 65.6 60.5 65.4 65.9 63.0 66.7 64.0 64.7
S 61.4 75.6 70.3 75.2 68.3 74.9 75.3 68.9 73.3 69.8 73.0
V 68.9 80.1 75.2 80.4 73.8 79.4 82.9 69.3 79.6 77.2 78.6
Table 12: Ablation study on hyperparameters for CLAP using ViT-B/16 on the VLCS dataset. We report domain-wise average top-1 accuracy (%) under three prompt settings—ZS(C), ZS(CP), and ZS(PC)—by varying: (1) $\tau$ (temperature), (2) latent dimension size, and (3) $\alpha$ (residual weight).
Hyper- Value Average Top-1 Accuracy (%) on VLCS (↑)
ZS(C) ZS(CP) ZS(PC)
C L S V C L S V C L S V
$\tau$ 0.1 99.9 67.6 77.5 84.2 99.9 70.9 74.9 85.9 99.9 71.2 74.6 86.3
0.3 99.9 66.3 77.2 82.4 99.9 69.9 76.7 85.2 99.9 69.9 76.4 85.4
0.5 99.9 67.7 78.0 84.9 99.9 70.4 77.2 86.0 99.9 70.7 76.9 86.2
0.7 99.9 65.9 77.7 83.1 99.9 68.9 77.9 84.9 99.9 69.6 77.7 85.0
0.9 99.9 66.0 77.6 83.3 99.9 69.0 77.9 85.0 99.9 69.7 77.5 85.0
Latent Dim. 128 99.9 66.0 77.6 82.6 99.9 70.0 77.4 85.4 99.9 70.1 77.1 85.7
192 99.9 64.9 77.9 83.0 99.9 68.9 78.0 85.6 99.9 69.0 77.8 86.0
256 99.9 63.8 77.6 82.7 99.9 67.6 78.7 84.8 99.9 67.8 78.6 85.2
320 99.9 66.0 77.8 82.9 99.9 69.2 78.1 85.3 99.9 69.7 77.7 85.5
384 99.9 65.8 76.9 82.8 99.9 69.4 77.5 85.3 99.9 69.6 77.0 85.5
448 99.9 65.8 77.4 82.1 99.9 69.7 77.6 84.9 99.9 69.9 77.1 85.6
512 99.9 67.7 78.0 84.9 99.9 70.4 77.2 86.0 99.9 70.7 76.9 86.2
$\alpha$ $10^{-1.5}$ 99.9 66.5 77.9 83.1 99.9 70.4 77.1 86.0 99.9 70.3 76.6 86.1
$10^{-1}$ 99.9 69.5 77.5 85.7 99.9 70.4 77.1 86.2 99.9 70.9 76.5 86.1
$10^{-0.5}$ 99.9 70.6 75.2 85.5 99.9 70.7 75.7 85.9 99.9 71.0 75.1 85.7
$10^{0}$ 99.8 71.5 73.5 83.5 99.9 71.7 74.4 85.8 99.8 72.3 73.5 85.5
$10^{0.5}$ 99.8 72.0 73.1 85.5 99.8 72.2 73.7 85.7 99.8 72.5 72.9 85.6
$10^{1}$ 99.8 72.1 72.8 85.4 99.8 72.3 73.4 85.7 99.8 72.5 72.9 85.5
$10^{1.5}$ 99.8 72.1 72.8 85.4 99.8 72.2 73.3 85.7 99.8 72.6 72.7 85.5
B.2 LLM Prompt Generation
We employ ChatGPT-3.5 [3] to generate the LLM-based prompts used in our comparative analysis of prompt sources. The procedure, illustrated in the following toolbox, involves prompting ChatGPT-3.5 to produce semantically rich, stylistically varied textual descriptions for specific class names. For each class, we generate 120 distinct samples. table 13 shows representative examples of the generated prompts.
LLM Prompting Procedure (ChatGPT-3.5) User: I have some class names, such as "dog", "car", "person", "chair". I want you to generate various styles of these objects in different art-style pictures. Provide me with a list of different sentences that describe a scene containing the object, with different object attributes (such as object size, condition or age (if applicable), colors), different backgrounds the object may appear in; different image types (such as photo, painting, etc). ChatGPT: Absolutely, I’ll provide you with a list of descriptive sentences for each of the class names "dog", "car", "person", "chair", incorporating various object attributes, backgrounds, and image types. Each sentence will describe a unique scene:
Table 13: Representative examples of LLM-generated prompts for selected classes. Each description was generated by ChatGPT-3.5 to simulate diverse visual and stylistic contexts.
Class Generated Prompt
Bird A pair of vibrant macaws converse in a lush, tropical rainforest, depicted in a lively, exotic wildlife painting.
A solitary eagle watches over a vast, rugged canyon at sunrise, portrayed in a majestic, wilderness landscape photograph.
Dog A sleek Whippet races in a competitive dog track, illustrated in a fast-paced, dynamic sports style.
A sturdy and reliable English Bulldog watching over a small shop, its solid presence reassuring to the owner.
Car A quirky art car parades through the streets in a colorful festival, captured in a fun, expressive style illustration.
A high-tech, autonomous car maneuvers through a smart city environment, portrayed in a futuristic, sci-fi digital art piece.
Chair A folding chair at an outdoor wedding, elegantly decorated and part of a beautiful ceremony.
A high-end executive chair in a law firm, projecting authority and professionalism.
Person An energetic coach motivates a team on a sports field, illustrated in an inspiring, leadership-focused painting.
A graceful figure skater glides across an ice rink, captured in a delicate, winter-themed pastel drawing.
B.3 Experiments on ViT-L/14
We refine the output dimension to match the ViT-L/14 input dimension of 768. The latent dimensions are set to 448 for PACS, 640 for VLCS, and 768 for both OfficeHome and DomainNet. The inference weighting parameter $\alpha$ is configured as 0.1 for PACS, 0.03 for VLCS, 0.14 for OfficeHome, and 0.2 for DomainNet. All other training configurations remain consistent with those used in the ViT-B/16 experiments. The training setup for Im.Aug mirrors that of CLAP for each dataset, with $\alpha$ set to 0.1 for PACS and 0.03 for the other three datasets.
Table 14 presents the zero-shot results for the ViT-L/14 model across four distinct prompts, following the same evaluation protocol as in the ViT-B/16 experiments. These results demonstrate that CLAP is more effective than Im.Aug in enhancing zero-shot performance. Furthermore, table 15 shows that CLAP substantially reduces performance variance across prompts, further validating its improvements over CLIP across multiple model scales. Detailed domain-level results are reported in table 16, providing a comprehensive breakdown of performance by domain.
B.4 Experiments on ResNet50x16
To validate our approach on different model structures, we repeated zero-shot experiments on the ResNet50x16 model pre-trained with CLIP. Since the output dimension of CLIP is the same as ViT-B/16, we used the same training configuration as ViT-B/16 for training Im.Aug and CLAP. For inference, we refined the weighting coefficient $\alpha$ to 0.1, 1, 0.03, and 0.1 for Im.Aug, and 0.03, 0.2, 0.06, and 0.1 for CLAP, for PACS, VLCS, OfficeHome, and DomainNet respectively.
Table 17 showcases the zero-shot results for ResNet50x16 model across different prompts, substantiating that CLAP is more effective than Im.Aug in refining CLIP features. Moreover, table 18 illustrates that both Im.Aug and CLAP reduce variations in zero-shot performance across different prompts, with the improvement of CLAP being more significant. The results validate our approach across different model scales, including both ViT-based and CNN-based structures. Domain-level results are detailed in table 19.
Appendix C Discussion
C.1 Rationale behind CLAP’s Foundation on CLIP
The primary challenge in cross-modal transferability lies in the significant domain gap between text and image data, which typically hinders the direct application of models trained in one modality to another. For a causal explaination, despite the consistency of the content variable that dictates the object label across modalities, the generative processes from latent variables to observations inherent to each modality differ markedly. The CLIP model, trained on a comprehensive dataset of image-text pairs with a symmetric InfoNCE loss, significantly ameliorates this issue. By aligning the features of text and images into similar patterns, it facilitates leveraging a network trained atop the CLIP encoder of one modality as a viable proxy for the other. Consequently, this allows for the direct application of the disentangled network trained in the text modality atop CLIP’s image encoder to refine representations.
C.2 Impact of Image and Text Augmentations
Identifying pure content factors poses a significant challenge. This difficulty primarily arises from the need for finding effective augmentations of observational data to alter style factors significantly while preserving content integrity.
Through the cross-modal alignment provided by CLIP, we discovered that disentangling in one modality can seamlessly improve representations in both modalities. The impact of image augmentations has been well-explored and found effective at preserving content, but traditional methods do not impose sufficient changes to remove all style information. Our exploration of text augmentations reveals that the logical structure of text and the relative ease of implementing style changes can have a significant impact on achieving disentanglement. However, more efficient methods are worthy of exploration.
A promising direction for future research is to explore efficient combinations of both modalities to enhance disentangled semantics. As each modality has its unique advantages—Text data recapitulates properties well since it is pre-processed by human intelligence, while image data is more precise in depicting the exact same objects or events due to its more detailed nature—the impact of combining augmentations of both modalities could be substantial.
Table 14: Zero-shot performance across four prompts ("C", "PC", "CP") and 1 noised prompts ("NC") with CLIP pre-trained ViT-L/14 model. CLAP demonstrates consistent gains in zero-shot performance across all datasets, validating its effectiveness.
Prompt Method Zero-shot performance, avg. top-1 acc. (%) (↑)
       PACS      VLCS OfficeHome DomainNet Overall
ZS(C) CLIP 97.6 77.1 85.9 63.2 80.9
Im.Aug 98.3 78.5 86.0 63.4 81.6
CLAP 98.5 80.7 87.5 64.2 82.7
ZS(CP) CLIP 97.3 80.6 86.0 62.0 81.5
Im.Aug 98.3 81.1 86.1 62.4 82.0
CLAP 98.5 81.4 87.9 63.7 82.9
ZS(PC) CLIP 98.4 81.7 86.5 63.5 82.5
Im.Aug 98.6 81.9 86.6 63.7 82.7
CLAP 98.6 82.2 88.0 64.5 83.3
ZS(NC) CLIP 91.0 65.5 77.1 55.4 72.3
Im.Aug 95.6 69.3 77.1 55.7 74.4
CLAP 98.5 73.1 81.3 58.3 77.8
Table 15: Prompt robustness analysis with CLIP pre-trained ViT-L/14 model. CLAP reduces the variance in zero-shot performance across different prompts.
Metric Method Zero-shot variance, avg. top-1 acc. (%) (↓)
       PACS      VLCS OfficeHome DomainNet Overall
$R$ CLIP 1.0 4.6 0.6 1.5 1.9
Im.Aug 0.3 3.4 0.6 1.3 1.4
CLAP 0.1 1.5 0.4 0.7 0.7
$\delta$ CLIP 0.4 2.0 0.3 0.6 0.8
Im.Aug 0.1 1.5 0.3 0.5 0.6
CLAP 0.0 0.6 0.2 0.3 0.3
$\Delta_{(NC)}$ CLIP 6.6 11.5 8.8 7.8 8.7
Im.Aug 2.7 9.2 8.9 7.7 7.1
CLAP 0.1 7.7 6.3 5.9 5.0
Table 16: Domain-level zero-shot results for the ViT-L/14 model across test datasets. CLAP consistently improves performance across all domains and prompt types, demonstrating its robustness and generalization capabilities.
Datasets Domains Domain-level avg. top-1 acc. (%) of zero-shot performance using ViT-L/14 (↑)
ZS(C) ZS(CP) ZS(PC) ZS(NC)
CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP
PACS A. 97.2 98.0 98.8 96.8 98.0 98.5 98.7 98.8 98.9 85.6 91.6 98.6
C. 99.5 99.6 99.8 98.3 99.6 99.7 99.5 99.6 99.7 95.9 98.1 99.6
P. 99.9 100.0 100.0 99.4 99.5 100.0 99.9 100.0 99.9 91.1 97.5 99.9
S. 93.8 95.7 95.5 94.8 96.0 95.7 95.4 95.9 95.8 91.5 95.2 95.8
VLCS C. 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 87.5 87.9 94.4
L. 57.4 60.1 64.3 71.3 71.6 72.6 71.7 72.0 72.6 53.8 59.7 60.7
S. 71.0 72.4 74.4 66.2 67.4 66.8 69.9 70.4 69.9 55.9 60.5 62.9
V. 80.0 81.6 84.3 85.2 85.7 86.2 85.1 85.3 86.4 65.0 69.3 74.3
O.H. A. 86.2 86.3 87.7 85.7 86.2 88.1 87.0 87.0 87.8 78.1 77.1 80.7
C. 73.3 73.4 75.7 73.8 73.4 76.0 73.1 73.5 76.0 65.9 66.3 70.6
P. 92.0 91.8 93.6 92.3 92.4 94.3 92.9 92.8 94.1 80.7 81.0 86.8
R. 92.2 92.7 93.0 92.2 92.4 93.4 93.1 93.3 93.9 83.8 84.0 86.9
D.N. C. 78.4 78.5 79.1 77.5 77.7 78.8 79.4 79.4 79.7 70.0 70.4 72.8
I. 52.9 53.0 54.6 50.4 50.7 53.6 51.7 52.0 53.9 45.3 45.2 48.8
P. 70.4 70.8 72.4 68.9 69.9 72.1 69.9 70.6 72.7 59.9 60.3 64.8
Q. 21.5 21.6 22.5 20.6 20.9 21.7 22.6 22.8 22.9 17.9 18.4 20.2
R. 85.8 85.9 85.9 85.3 85.5 85.7 86.3 86.4 86.2 77.5 77.5 78.7
S. 70.2 70.4 70.7 69.4 69.8 70.6 71.0 71.3 71.5 62.0 62.2 64.6
Table 17: Zero-shot performance with the CLIP pre-trained ResNet50x16 model. CLAP consistently enhances performance across all datasets, validating its effectiveness.
Prompt Method Zero-shot performance, avg. top-1 acc. (%) (↑)
       PACS      VLCS OfficeHome DomainNet Overall
ZS(C) CLIP 96.1 70.4 80.4 57.1 76.0
Im.Aug 96.4 74.7 80.4 57.1 77.2
CLAP 97.0 79.9 81.6 58.0 79.1
ZS(CP) CLIP 95.0 73.5 79.0 56.1 75.9
Im.Aug 95.7 75.8 79.3 56.5 76.8
CLAP 96.7 80.3 79.9 57.4 78.6
ZS(PC) CLIP 96.5 78.4 81.7 57.1 78.4
Im.Aug 97.0 79.8 81.8 57.4 79.0
CLAP 96.8 80.1 82.5 58.2 79.4
ZS(NC) CLIP 86.4 61.2 69.3 48.2 66.3
Im.Aug 88.3 71.3 69.5 48.7 69.4
CLAP 94.9 80.1 71.9 50.6 74.4
Table 18: CLAP consistently reduces variance in zero-shot performance across different prompts using the CLIP pre-trained ResNet50x16 model, further validating its robustness and effectiveness.
Metric Method Zero-shot variance, avg. top-1 acc. (%) (↓)
       PACS      VLCS OfficeHome DomainNet Overall
$R$ CLIP 1.5 8.0 2.7 1.1 3.3
Im.Aug 1.3 5.1 2.5 0.9 2.4
CLAP 0.3 0.4 2.6 0.8 1.0
$\delta$ CLIP 0.6 3.3 1.1 0.5 1.4
Im.Aug 0.5 2.2 1.0 0.4 1.0
CLAP 0.1 0.2 1.1 0.3 0.4
$\Delta_{(NC)}$ CLIP 9.7 9.3 11.1 8.9 9.7
Im.Aug 8.1 3.5 10.9 8.5 7.7
CLAP 2.1 -0.1 9.7 7.5 4.8
Table 19: Domain-level zero-shot performance using ResNet50x16 across test datasets. CLAP consistently improves upon CLIP and Im.Aug across domains and prompt types (ZS(C), ZS(CP), ZS(PC), ZS(NC)), demonstrating its generalizability and robustness.
Datasets Domains Domain-level avg. top-1 acc. (%) of zero-shot performance using RN50x16 (↑)
ZS(C) ZS(CP) ZS(PC) ZS(NC)
CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP
PACS A. 95.7 95.8 97.2 93.7 95.5 96.5 95.7 96.7 96.8 81.2 84.0 94.5
C. 98.3 98.2 99.0 98.1 98.8 99.0 98.6 98.7 98.9 92.3 93.2 98.0
P. 98.9 98.6 99.9 98.4 97.8 99.8 99.8 99.9 99.9 85.3 87.3 95.2
S. 91.5 93.1 91.9 89.7 90.8 91.5 91.8 92.9 91.6 86.9 88.6 92.1
VLCS C. 96.8 97.1 99.3 99.7 99.4 99.3 99.7 99.6 99.4 75.6 89.3 99.4
L. 53.4 60.8 65.9 51.6 58.9 67.3 59.5 68.1 66.8 54.1 60.6 67.0
S. 63.2 70.9 69.5 68.0 72.9 69.5 72.9 73.7 69.0 52.0 66.7 69.5
V. 68.4 70.1 85.2 74.5 72.1 85.3 81.7 78.0 85.2 63.1 68.5 84.5
O.H. A. 82.2 82.5 83.5 79.7 79.9 80.6 82.0 82.4 83.4 67.7 68.9 72.1
C. 63.0 62.9 64.7 61.7 62.2 62.8 65.4 65.3 66.1 54.6 55.0 56.8
P. 88.2 87.9 89.0 87.4 87.5 88.5 90.0 89.9 90.6 75.4 75.3 78.2
R. 88.1 88.2 89.1 87.3 87.5 87.6 89.2 89.5 89.7 79.5 78.9 80.3
D.N. C. 69.0 68.9 69.6 68.6 68.6 69.4 69.9 70.0 70.4 59.5 60.1 61.4
I. 51.0 51.1 52.7 48.2 49.0 50.6 48.2 48.9 50.7 41.2 41.6 44.3
P. 65.2 65.6 66.5 63.7 64.4 65.6 65.4 65.9 67.0 53.5 54.3 56.8
Q. 11.8 11.9 12.7 12.3 12.6 13.1 11.8 12.2 12.7 9.3 9.7 11.0
R. 82.1 82.2 83.1 81.6 81.8 82.6 83.3 83.4 83.8 72.9 73.0 74.7
S. 63.2 63.1 63.6 62.0 62.4 63.4 63.9 63.8 64.6 53.1 53.4 55.3

### v6

@@
 Abstract
-Contrastive vision-language models, such as CLIP, have garnered considerable attention for various downstream tasks, mainly due to the remarkable generalization ability of the learned features.
-However, the features they learn often blend content and style information, which somewhat limits their generalization capabilities under distribution shifts.
+Contrastive vision-language models, such as CLIP, have garnered considerable attention for various downstream tasks, mainly due to the remarkable ability of the learned features for generalization.
+However, the features they learned often blend content and style information, which somewhat limits their generalization capabilities under distribution shifts.
 To address this limitation, we adopt a causal generative perspective for multimodal data and propose contrastive learning with data augmentation to disentangle content features from the original representations.
-To achieve this, we begin by exploring image augmentation techniques and develop a method to seamlessly integrate them into pre-trained CLIP-like models to extract pure content features.
-Taking a step further, and recognizing the inherent semantic richness and logical structure of text data, we explore the use of text augmentation to isolate latent content from style features.
-This enables CLIP-like models’ encoders to concentrate on latent content information, refining the representations learned by pre-trained CLIP-like models.
+To achieve this, we begin with exploring image augmentation techniques and develop a method to seamlessly integrate them into pre-trained CLIP-like models to extract pure content features.
+Taking a step further, recognizing the inherent semantic richness and logical structure of text data, we explore the use of text augmentation to isolate latent content from style features.
+This enables CLIP-like model’s encoders to concentrate on latent content information, refining the learned representations by pre-trained CLIP-like models.
 Our extensive experiments across diverse datasets demonstrate significant improvements in zero-shot and few-shot classification tasks, alongside enhanced robustness to various perturbations.
-These results underscore the effectiveness of our proposed methods in refining vision-language representations and advancing the state of the art in multimodal learning.
-Keywords Data Augmentation  $\cdot$ Latent Variables  $\cdot$ Disentanglement
+These results underscore the effectiveness of our proposed methods in refining vision-language representations and advancing the state-of-the-art in multimodal learning.
+[^1] Our code is available at https://github.com/YichaoCai1/CLAP.
+Keywords:
+Data Augmentation Latent Variables Disentanglement
 1 Introduction
-Vision-language models, exemplified by CLIP [36], have garnered substantial attention due to their exceptional generalization capabilities, achieved through the learned features obtained by utilizing a cross-modal contrastive loss [36, 20, 25].
+Vision-language models, exemplified by CLIP [36], have garnered substantial attention due to their exceptional generalization capabilities, achieved through the learned features, obtained by utilizing a cross-modal contrastive loss [36, 20, 25].
 However, despite being pre-trained on extensive datasets, CLIP-like models fall short in disentangling latent content information and latent style information.
 Consequently, they are not immune to spurious correlations, i.e., style-related information is erroneously utilized to predict task-related labels.
-These limitations become evident in the presence of distribution shifts or adversarial attacks, where spurious correlations often change across different environments.
-For example: (1) a notable dependence on specific input text prompts has been reported for zero-shot capabilities [48, 47, 21]; (2) performance decline in few-shot scenarios has been observed [36, 13]; and (3) susceptibility to adversarial attacks has been explored [33, 45, 43].
+These limitations become evident in the presence of distribution shifts or adversarial attacks where spurious correlations often change across different environments.
+For examples, (1) a notable dependence on specific input text prompts has been reported for zero-shot capabilities [48, 47, 21]; (2) performance decline in few-shot scenarios has been observed in few-shot learning scenarios [36, 13]; and (3) susceptibility to adversarial attacks has been explored [33, 45, 43].
 […]
@@
 (c) A qualitative comparison of image features for zero-shot classification using "a photo of a [class]" prompts, visualized using class activation map (CAM) [32], demonstrates that while image augmentation can enhance CLIP features, the features obtained through text augmentation methods predominantly focus on the content.
-Taking a causal perspective, this work begins with a simple yet effective method—image augmentation—to disentangle content and style information within the learned representations of CLIP-like models.
-This approach is inspired by recent advancements in theoretical developments in causal representation learning [41], which demonstrate that augmented images can be interpreted as results of soft interventions on latent style variables, as depicted in fig.
+Taking a causal perspective, this work begin with a simple yet effective method, image augmentation, to disentangle content and style information within the learned representations of CLIP-like models.
+This approach is inspired by recent advancements in theoretical development in causal representation learning [41], which demonstrate that augmented image can be interpreted as a result of soft interventions on latent style variables, as depicted in Fig.
 1.
@@
 Consequently, using contrastive learning, it becomes feasible to isolate the invariant content information from the variant style information.
-Motivated by this theoretical advancement, we propose a practical method to incorporate image augmentation into CLIP-like models to extract content information from the originally learned features.
-Specifically, a disentangled network is designed to fine-tune the pre-trained CLIP model using contrastive loss with image augmentation.
-Despite the advancements made in disentangling content and style information from the original features learned by CLIP-like models through image augmentation, we recognize an inherent limitation: it is generally challenging to design adequate image augmentations that ensure all style factors change in an image.
+Motivated by this theoretical advancement, we propose a practical method to incorporate image augmentation into CLIP-like models to extract content information from the original learned features.
+Specifically, a disentangled network is designed to fine-tune the pre-trained CLIP model by using a contrastive loss with image augmentation.
+Despite the advancements made in disentangling content and style information from the original features learned by CLIP-like models through image augmentation, we recognize an inherent limitation: it is generally challenging to design adequate image augmentations to ensure all style factors change in an image.
 Theoretically, disentangling content and style information necessitates changes in all style factors [41].
 However, inducing sufficient changes in latent style through image augmentation poses challenges due to the high dimensionality and complexity of style information in image data.
-Achieving significant style variation via artificially designed image augmentation techniques—such as transforming a photograph of a dog into a sketch while preserving content but dramatically altering style—is notably difficult.
+Achieving significant style variation via artificially designed image augmentation techniques, such as transforming a photograph of a dog into a sketch while preserving content but dramatically altering style, is notably difficult.
 Taking a further step, rather than relying on image augmentation, we explore the use of text augmentation to disentangle latent content and style factors.
-This shift is motivated by two key observations: (1) vision and language data share the same latent space, and therefore, text augmentation can also be utilized to induce changes in latent style factors; (2) text data inherently possesses high semanticity and logical structure, making it more amenable to property-wise manipulation compared to image data.
-Consequently, implementing sufficient style changes through text augmentation is more feasible than image augmentation, contributing to isolating content from style information (see fig.
-1 for visual comparison).
+This shift is motivated by two key observations: 1) Vision and language data share the same latent space.
+Therefore, text augmentation can also be utilized to induce changes in latent style factors instead of image augmentation.
+2) Text data inherently possesses high semanticity and logical structure, making it more amenable to property-wise manipulation compared to image data.
+Consequently, implementing sufficient style changes through text augmentation is more feasible than image augmentation, contributing to isolating content from style information, see Fig.
+1 for visual comparison.
 For instance, transforming text from "a photo of a dog" to "a sketch of a dog" is straightforward in the language modality, whereas achieving a similar transformation in image data is challenging.
-Inspired by these observations, we posit that introducing style variations through text augmentation, as illustrated in fig.
-1, provides a more effective approach for learning vision-language content features than relying solely on image augmentation.
-In summary, our contributions include:
-  1. 1.
-    Aimed at disentangling latent content and style factors to refine vision-language features of pre-trained CLIP-like models, we propose contrastive learning with data augmentation to fine-tune the original features from a causal perspective.
-  2. 2.
-    We present a novel method customized for pre-trained CLIP-like models.
-This method leverages a disentangled network, which is trained using contrastive learning with image augmentation, to extract latent content features from the learned features provided by the image encoder.
-  3. 3.
-    We propose Contrastive Learning with Augmented Prompts (CLAP) to extract latent content features from representations of CLIP-like models.
-It begins by training a disentangled network using the pre-trained text encoder and text augmentation.
-Subsequently, the trained disentangled network is transferred to the image encoder.
-  4. 4.
-    Experiments conducted on a large real-world dataset demonstrate the effectiveness of the proposed image and text augmentations in improving zero-shot and few-shot performance, as well as robustness against perturbations.
+Inspired by these observations, we posit that introducing style variations through text augmentation, as illustrated Fig.
+1, provides a more effective approach for learning vision-language content features than relying on image augmentation.
+In summary, our contributions include: (1) Aimed at disentangling latent content and style factors to refine vision-language features of pre-trained CLIP-like models, we propose constrastive learning with data augmentation to fine tune the original features of pre-trained CLIP-like models from a causal perspective.
+(2) We present a novel method customized for pre-trained CLIP-like models.
+This method leverages a disentangled network, which is trained using contrastive learning with image augmentation, to extract latent content features from the learned features provided by image encoder of CLIP-like models.
+(3) We propose Contrastive Learning with Augmented Prompts (CLAP), to extract latent content features from representations of CLIP-like models.
+It begins by training a disentangled network using the pre-trained text encoder of CLIP-like models and text augmentation.
+Subsequently, the trained disentangled network is transferred to the image encoder of CLIP-like models.
+(4) Experiments conducted on a large real dataset demonstrate the effectiveness of the proposed image augmentation and text augmentation in terms of zero-shot and few-shot performance, as well as robustness against perturbations.
 2 Related Work
-2.1 Contrastive Vision-Language Models
-Using a cross-modal contrastive loss, CLIP [36] introduced a scalable contrastive vision-language model by leveraging a large corpus of internet-sourced image-text pairs, demonstrating unprecedented zero-shot learning capabilities and exceptional generalization ability across datasets, while supporting numerous downstream tasks [38].
-ALIGN [20] expanded the scale of contrastive vision-language modeling by training on up to one billion image-text pairs and integrating the vision transformer’s self-attention mechanism [11], which further enhanced performance.
+2.0.1 Contrastive Vision-Language Models
+Using a cross-modal contrastive loss, CLIP [36] revolutionarily introduced a scalable contrastive vision-language model by leveraging a large corpus of internet-sourced image-text pairs, demonstrating unprecedented zero-shot learning capabilities and exceptional generalization ability across datasets and supporting numerous downstream tasks [38].
+ALIGN [20] expanded the scale of contrastive vision-language modeling, training on up to one billion image-text pairs while integrating the vision transformer’s self-attention mechanism [11], which further enhanced performance.
 Despite their successes, CLIP-like models exhibit sensitivity to input text prompts [48, 21], leading to variable performance across different prompts.
@@
 Furthermore, CLIP-like models are vulnerable to adversarial attacks [4, 12], with current strategies [33, 45] involving adversarial-natural image pairs to improve resilience.
-Our work diverges from these task-specific approaches by aiming to enhance CLIP’s representations from a disentanglement perspective, addressing the aforementioned issues inherent in CLIP-like models.
-2.2 Disentangled Representation Learning
+Our work diverges from task-specific approaches by aiming to enhance CLIP’s representations from a disentanglement perspective, addressing the aforementioned issues inherent in CLIP-like models.
+2.0.2 Disentangled Representation Learning
 Aimed at segregating intrinsic latent factors in data into distinct, controllable representations, disentangled representation learning benefits various applications [44, 40, 24].
-Specifically, in classification tasks, it has been shown that improving model performance and robustness against data distribution perturbations can be achieved by more effectively disentangling invariant content variables, without needing to fully identify all intrinsic latent variables [22, 28, 26, 27].
-Within single modalities, studies such as [49] have illustrated that contrastive learning [7, 16, 18] can potentially reverse the data-generative process, aiding in the separation of representations.
+Specifically, in classification tasks, it’s shown that enhancing the model’s performance and robustness against data distribution perturbations can be achieved by more effectively disentangling invariant content variables, without needing to identify all intrinsic latent variables completely [22, 26, 27, 28].
+Within single modalities, studies such as [49] have illustrated that contrastive learning [7, 16, 18] can potentially reverse the data generative process, aiding in the separation of representations.
 Furthermore, [41] suggest that image augmentation can help isolate content variables from the latent space through significant stylistic changes.
 [19] employs mixture techniques for data augmentation, enabling more abundant cross-modal matches.
-Diverging from these methods, our approach focuses on employing text augmentation to disentangle latent content variables, introducing a unique approach to learning refined vision-language representations.
-3 A Causal Generative Model for Multimodal Data
-To understand pre-trained CLIP-like models, we investigate the underlying causal generative process for vision-language data.
-We consider the following causal generative model, as depicted in fig.
+Diverging from these methods, our approach focuses on employing text augmentation to disentangle latent content variables, introducing a unique approach to learn refined vision-language representations.
+3 A Causal Generative Model for Multi-Modal Data
+To understand pretrained CLIP-like models, we investigate the underlying causal generative process for vision-language data.
+We consider the following causal generative model as depicted in Fig.
 1.
-In the proposed model, the shared latent space governing vision and language data is divided into two distinct subspaces: one corresponding to the latent content variables $\mathbf{c}$ and the other to the latent style variables $\mathbf{s}$.
+In the proposed model, the shared latent space ruling vision and language data is divided into two distinct sub-spaces: one corresponding to the latent content variables $\mathbf{c}$ and the other to the latent style variables $\mathbf{s}$.
 The latent content variables are posited to determine the object label $\mathbf{y}$, a relationship corroborated by prior studies [22, 29, 31].
 Furthermore, to elucidate the correlation between the latent style variable $\mathbf{s}$ and the object variable $\mathbf{y}$, our model incorporates the premise that the latent content variable $\mathbf{c}$ causally influences the latent style variable $\mathbf{s}$, in concordance with the principles of causal representation learning highlighted in recent literature [10, 29, 41].
-Additionally, considering the diversity between image data and text data—where information in image data is typically much more detailed, while information in text data tends to be more logically structured—we posit distinct causal mechanisms for the respective generative processes.
-Our causal generative model is formulated using structural causal models [2]:
-$\mathbf{s}:=\mathbf{g}_{s}(\mathbf{c}),~{}\mathbf{x}:=\mathbf{g}_{x}(\mathbf{c,s}),~{}\mathbf{t}:=\mathbf{g}_{t}(\mathbf{c,s}),~{}\mathbf{y}:=\mathbf{g}_{y}(\mathbf{c}).$ (1)
+Additionally, considering the diversity between image data and text data, where information in image data is typically much more details while information in text data tends to be more logically structured nature, we posit distinct causal mechanisms for the generation processes.
+Our causal generative model is formulated as following structural causal models [2]:
+$\mathbf{s}:=\mathbf{g_{s}(c)},\quad\mathbf{x}:=\mathbf{g_{x}(c,s)},\quad\mathbf{t}:=\mathbf{g_{t}(c,s)},\quad\mathbf{y}:=\mathbf{g_{y}(c)}.$ (1)
 In Eq.
 1, the style variable $\mathbf{s}$ is causally influenced by the content via $\mathbf{g_{s}}$; $\mathbf{x}$ and $\mathbf{t}$ denote visual and textual data, respectively.
-Both visual and textual data are causally produced by the shared latent variables $\mathbf{c}$ and $\mathbf{s}$ through distinct, reversible generative processes: $\mathbf{g}_{x}$ for images and $\mathbf{g}_{t}$ for text.
-The label $\mathbf{y}$ of a sample is exclusively determined by the content variable $\mathbf{c}$ via $\mathbf{g}_{y}$.
-For simplicity, exogenous noise variables are implicitly assumed but not explicitly represented in the formulation of the causal generative model, in line with the common understanding that each latent variable is influenced by exogenous noise.
-Recent seminal work in [41] has demonstrated that the latent content variable $\mathbf{c}$ can be identified up to block identifiability (i.e., $\mathbf{c}$ can be isolated from the style variable $\mathbf{s}$) by requiring all latent style variables to change (e.g., soft interventions on all latent style variables).
-This change can be achieved through image augmentation; that is, the augmented image $\mathbf{\tilde{x}}$ can be interpreted as a generative result of $\mathbf{\tilde{s}}$, which is produced through soft interventions on the original latent style variables $\mathbf{s}$.
-Despite this theoretical advancement, the practical implementation of this result within CLIP-like models remains unclear.
-In this study, we propose a practical method to disentangle content and style information within CLIP-like models by employing image augmentation, as detailed in § 4.1.
-Moreover, we recognize that implementing sufficient changes in all latent style variables $\mathbf{s}$ through text augmentation is more feasible than through image augmentation.
-Due to the high semanticity and logical structure of text data, we further explore the use of text augmentation to separate content information from style information, as discussed in § 4.2.
+Both visual and textual data are causally produced by the shared latent variables $\mathbf{c}$ and $\mathbf{s}$ through distinct, reversible generative processes: $\mathbf{g_{x}}$ for images and $\mathbf{g_{t}}$ for text data, respectively.
+The label $\mathbf{y}$ of a sample is exclusively determined by the content variable $\mathbf{c}$ via $\mathbf{g_{y}}$.For simplicity, exogenous noises are implicitly assumed but not explicitly represented in the causal generative model’s formulation, aligning with the common understanding that each latent variable is influenced by exogenous noise.
+Recent seminal work in [41] has demonstrated that the latent content variable $\mathbf{c}$ can be identified up to block identifiability (i.e., $\mathbf{c}$ can be isolated from style variable $\mathbf{s}$), by requiring all latent style variables to change (e.g., soft interventions on all latent style variables).
+This change can be achieved through image augmentation, i.e., the augmented image $\mathbf{\tilde{x}}$ can be interpreted as a generative result of $\mathbf{\tilde{s}}$, which is produced through soft interventions on original latent style variables $\mathbf{s}$.
+Despite such theoretical advancement, the practical implementation of this theoretical result within CLIP-like models remains unclear.
+In this study, we propose a practical method to disentangle content and style information within CLIP-like models by employing image augmentation, as detailed in Sec.
+4.1.
+Moreover, we recognize that implementing sufficient changes on all latent style variables $\mathbf{s}$ through text augmentation is more feasible than image augmentation, due to high semanticity and logical structure in text data, we delve into the use of text augmentation to separate content information from style information, as discussed in Sec.
+4.2.
 4 Isolating Content from Style with Data Augmentation
-In this section, we propose the use of data augmentation to extract content information from the learned features in pre-trained CLIP-like models.
+In this section, we propose the employment of data augmentation to extract content information from the learned features in pre-trained CLIP-like models.
 Essentially, data augmentation facilitates the alteration of style factors while preserving content factors.
 Consequently, leveraging contrastive learning enables the segregation of content information from style information.
-We explore two distinct forms of data augmentation, namely image augmentation (§ 4.1) and text augmentation (§ 4.2).
+We delve into two distinct forms of data augmentation, namely image augmentation (Sec.
+4.1) and text augmentation (Sec.
+4.2).
 4.1 Isolating Content from Style with Augmented Images
-While recent studies (von et al., 2021) have provided theoretical assurance regarding the disentanglement of content and style through contrastive learning with data augmentation, it remains unclear how these findings can be applied to vision-language models.
-We extend these theoretical results to CLIP-like models as follows.
-The theoretical framework suggests using InfoNCE loss [34] to extract content information, as formulated below:
-$\mathcal{L}(\mathbf{f};\{\mathbf{x}_{i},\mathbf{\tilde{x}}_{i}\}_{i=1}^{b},\tau)=-\frac{1}{b}\sum\nolimits_{i=1}^{b}\log\frac{\exp\bigl{(}\langle\mathbf{f}(\mathbf{x}_{i}),\mathbf{f}(\mathbf{\tilde{x}}_{i})\rangle/\tau\bigr{)}}{\sum_{j=1}^{b}\exp\bigl{(}\langle\mathbf{f}(\mathbf{x}_{i}),\mathbf{f}(\mathbf{\tilde{x}}_{j})\rangle/\tau\bigr{)}},$ (2)
-[…]
+While recent studies (von et al., 2021) have offered assurance regarding the disentanglement of content and style through contrastive learning with data augmentation, it remains unclear how these theoretical findings can be applied to the realm of vision-language models.
+We convert the theoretical findings into CLIP-like models in the following.
+The theoretical findings suggest using InfoNCE loss [34] to extract content information, as outlined below:
+$\mathcal{L}{(\mathbf{f};\{\mathbf{x}_{i},\mathbf{\tilde{x}}_{i}\}_{i=1}^{b},\tau)}=-\frac{1}{b}\sum\nolimits_{i=1}^{b}\mathrm{log}\frac{\mathrm{exp}\left[\langle\mathbf{f}(\mathbf{x}_{i}),\mathbf{f}(\mathbf{\tilde{x}}_{i})\rangle/\tau\right]}{\sum_{j=1}^{b}\mathrm{exp}\left[\langle\mathbf{f}(\mathbf{x}_{i}),\mathbf{f}(\mathbf{\tilde{x}}_{j})\rangle/\tau\right]},$ (2)
+where $\{\mathbf{x}_{i}\}_{i=1}^{b}$ represents a batch of $b$ samples from the training dataset, $\mathbf{f}(\mathbf{x}_{i})$ denotes sample $\mathbf{x}_{i}$’s features through model $\mathbf{f}$, $\mathbf{\tilde{x}}_{i}$ is the augmented counterpart of $\mathbf{x}_{i}$, and $\mathrm{\langle\mathbf{z}_{1},\mathbf{z}_{2}\rangle}$ represents the cosine similarity between two feature vectors, $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$, and $\tau$ represents the temperature parameter influencing the loss.
 (a) Image augmentation
@@
 Figure 2: Refining CLIP through data augmentation.
-(a) Training involves a disentangled network $\mathbf{f}_{c}$, utilizing contrastive loss on original and augmented image pairs $\mathbf{x}$ and $\mathbf{\tilde{x}}$, with CLIP’s image encoder $\mathbf{f}^{*}_{x}$ having frozen gradients.
+(a) Training involves a disentangled network $\mathbf{f_{c}}$, utilizing contrastive loss on original and augmented image pairs $\mathbf{x}$ and $\mathbf{\tilde{x}}$, with CLIP’s image encoder $\mathbf{f^{*}_{x}}$ holding frozen gradients.
 (b) More efficient content feature learning is achieved through contrastive learning with augmented text prompts $\mathbf{t}$ and $\mathbf{\tilde{t}}$, using the fixed text encoder $\mathbf{f^{*}_{t}}$ of CLIP.
 (c) Inference stage: The trained disentangled network $\mathbf{f^{*}_{c}}$ integrates with CLIP’s text and image encoders, $\mathbf{f^{*}_{t}}$ and $\mathbf{f^{*}_{x}}$, to enable zero-shot inference for an input image $\mathbf{x}$ and class names $\mathbf{t}_{1}$ to $\mathbf{t}_{n}$.
-We extend this formulation to refine pre-trained vision-language models using contrastive learning with augmented images (hereinafter referred to as "Im.Aug").
+We extend it to refine pre-trained vision-language models, utilizing contrastive learning with augmented images (hereinafter referred to as "Im.Aug").
 As illustrated in Fig.
 2(a), we train a disentangled network on top of CLIP’s pre-trained image encoder.
-To improve training efficiency and usability, we freeze the pre-trained image encoder.
-Based on the InfoNCE loss, the learning objective of Im.Aug is defined as:
-$\mathbf{f}^{*}_{c}=\mathop{\mathrm{argmin}}_{\mathbf{f}_{c}}\mathop{\mathbb{E}}_{\{\mathbf{x}_{i}\}_{i=1}^{b}\in\mathcal{D}_{x}}\mathcal{L}(\mathbf{f}_{c}\circ\mathbf{f}^{*}_{x};\{\mathbf{x}_{i},\mathbf{\tilde{x}}_{i}\}_{i=1}^{b},\tau),$ (3)
-where $\mathcal{D}_{x}$ denotes the training image dataset, $b$ is the batch size, and $\mathbf{f}_{c}$ is the disentangled network being trained.
-The pre-trained CLIP image encoder is represented by $\mathbf{f}^{*}_{x}$, with the asterisk "*" indicating that its weights are fixed.
-The variable $\mathbf{x}_{i}$ refers to an image sampled from $\mathcal{D}_{x}$, and $\mathbf{\tilde{x}}_{i}$ is its augmented view.
-The composition of the training dataset $\mathcal{D}_{x}$, the image augmentation techniques used, the structure of the disentangled network $\mathbf{f}_{c}$, and the usage of $\mathbf{f}^{*}_{c}$ after training are detailed in the following subsections.
-Data Synthesis and Image Augmentation
+To enhance training efficiency and the usability of the proposed method, we freeze the pre-trained image encoder.
+Based on an InfoNCE loss, the learning objective of Im.Aug is formulated as follows:
+$\mathbf{f^{*}_{c}}=\mathop{\mathrm{argmin}}_{\mathbf{f_{c}}}\mathop{\mathbb{E}}_{\{\mathbf{x}_{i}\}_{i=1}^{b}\in\mathcal{D}_{\mathbf{x}}}\mathcal{L}(\mathbf{f_{c}\circ f^{*}_{x}};\{\mathbf{x}_{i},\mathbf{\tilde{x}}_{i}\}_{i=1}^{b},\tau),$ (3)
+where $\mathcal{D}_{\mathbf{x}}$ denotes the training image dataset and $b$ represents the batch size, $\mathbf{f_{c}}$ is the disentangled network undergoing training.
+The pre-trained CLIP image encoder is represented by $\mathbf{f}^{*}_{\mathbf{x}}$, with the asterisk "*" signifying that the model weights remain fixed.
+The variable $\mathbf{x}_{i}$ refers to an image sampled from $\mathcal{D}_{\mathbf{x}}$, and $\mathbf{\tilde{x}}_{i}$ is its augmented view.
+The composition of the training dataset $\mathcal{D}_{\mathbf{x}}$, the image augmentation techniques used, the structure of the disentangled network $\mathbf{f_{c}}$, and the utilization of $\mathbf{f^{*}_{c}}$ post-training are detailed in the following subsections.
+4.1.1 Data Synthesis and Image Augmentation
 To generate training image data, we combine class names with various image and object attributes to create text prompts for each class.
-Using a stable diffusion model [39], we produce synthetic images that comprise our training dataset $\mathcal{D}_{x}$.
+Using a stable diffusion model [39], we produce synthetic images that comprise our training dataset $\mathcal{D}\mathbf{x}$.
 The creation of template prompts for stable diffusion is based on attributes such as object size, color, image type, and art style.
-As detailed in table 1, the attributes include 10 colors and 3 object sizes, as well as 8 image types and 2 art styles.
-By assembling these attributes into prompts such as "a [art style] [image type] of a [object size] [object color] [class]", we generate 480 unique texts per class, from which one image per prompt is synthesized.
-Further details on image synthesis and examples are provided in § B.1.
+As detailed in Tab.
+1, the attributes include 10 colors and 3 sizes for objects, and 8 types and 2 art styles for images.
+By assembling these attributes into prompts like "a [art style] [image type] of a [object size] [object color] [class]", we generate 480 unique texts for each class, from which one image per prompt is synthesized.
+Further details on image synthesis and examples are available in Sec.
+0.B.1.
 Table 1: Template-based prompts.
-Attributes used to generate text prompts follow the structured format: "a [art style] [image type] of a [object size] [object color] [class]", where "[class]" represents the class names.
+Attributes used to generate text prompts follow the structured format "a [art style] [image type] of a [object size] [object color] [class]", where "[class]" represents the class names.
 Object Color Object Size Image Type Art Style
-yellow, green, black, blue, multicolored, orange, red, white, brown, purple large, small, normal sized painting, cartoon, infograph, sketch, photograph, clipart, mosaic art, sculpture realistic, impressionistic
-For the image augmentation procedures, we adopt techniques commonly used in contrastive learning [7, 41, 8], specifically random cropping and color distortion.
-Disentangled Network Structure
+yellow, green, black, large,  painting, cartoon, infograph, realistic,
+blue, multicolored, orange, small, sketch, photograph, clipart,  impressionistic
+red, white, brown, purple  normal sized mosaic art, sculpture
+For the image augmentation procedures, we adopt techniques commonly used in contrastive learning practice [7, 41, 8], specifically random cropping and color distortion.
+4.1.2 Disentangled Network Structure
 Since the training process is based on CLIP’s pre-trained lower-dimensional features, our disentangled network adopts a multi-layer perceptron (MLP) architecture.
@@
 3.
-This design enables learning directly from the pre-trained representation space, avoiding a random starting point.
-It is inspired by ControlNet’s zero-conv operation [46], which we adapt to a zero-linear operation within our residual MLP.
-Within this architecture, the main branch includes a zero-initialized, bias-free linear layer, positioned after a SiLU activation and a normally initialized linear layer.
-Conventionally, the dimensions of features before the initial linear layer, between the first and second linear layers, and after the second linear layer are referred to as the input ($d_{\text{in}}$), latent ($d_{\text{mid}}$), and output ($d_{\text{out}}$) dimensions, respectively.
-To rectify any mismatches between the input and output dimensions, the network employs nearest-neighbor downsampling in the shortcut path, thereby ensuring both dimensional alignment and the preservation of sharpness in the input features.
-During inference, a weighting parameter $\alpha>0$ is introduced to modulate the contribution of features from the main branch before integration with the input features, whereas this parameter remains fixed at 1 during training.
-Inference
+This design enables learning directly from the pre-trained representation space, avoiding a random starting point, inspired by ControlNet’s zero-conv operation [46], which we adapt to a zero-linear operation within our residual MLP.
+Within this architecture, the main branch includes a zero-initialized, bias-free linear layer positioned subsequent to the combination of a SiLU activation and a normally initialized linear layer.
+Conventionally, the dimensions of features before the initial linear layer, situated between the first and second linear layers, and following the second linear layer, are named as the input $d_{in}$, latent $d_{mid}$, and output $d_{out}$ dimensions, respectively.
+To rectify any mismatches between the input and output dimensions, the network employs nearest-neighbor downsampling within the shortcut path, thereby ensuring both alignment and the preservation of sharpness for the input features.
+During the inference stage, a weighting parameter $\alpha>0$ is introduced to modulate the portion of features emanating from the main branch before their integration with the input features, whereas this parameter remains constant at 1 throughout the training phase.
+4.1.3 Inference
 After training, the disentangled network $\mathbf{f^{*}_{c}}$ is utilized following CLIP’s image encoder to extract visual content features.
-Moreover, given that vision-language data generation is rooted in a unified latent space, as described in § 3, $\mathbf{f}^{*}_{c}$ can be seamlessly integrated with CLIP’s image and text encoders to enhance zero-shot capabilities.
+Moreover, given that vision-language data generation is rooted in a unified latent space, as depicted in Sec.
+3, $\mathbf{f^{*}_{c}}$ can be seamlessly integrated with CLIP’s image and text encoders to enhance zero-shot capabilities.
 As shown in Fig.
-2(c), for an image $\mathbf{x}$, the operation is formulated as the composition function $\mathbf{f}^{*}_{c}\circ\mathbf{f}^{*}_{x}(\mathbf{x})$, and similarly, for a text $\mathbf{t}$, as $\mathbf{f}^{*}_{c}\circ\mathbf{f}^{*}_{t}(\mathbf{t})$.
-This integration preserves CLIP’s zero-shot functionality while providing refined features through improved disentanglement of content.
+2(c), for an image $\mathbf{x}$, the operation is formulated as the composition function $\mathbf{f^{*}_{c}\circ f^{*}_{x}(x)}$, and similarly, for a text $\mathbf{t}$, as $\mathbf{f^{*}_{c}\circ f^{*}_{t}(t)}$.
+This integration preserves CLIP’s zero-shot functionality while achieving refined features through the improved disentanglement of content.
 Figure 3: Structure of the disentangled network.
-The architecture encompasses a residual block featuring a zero-initialized, bias-free linear layer to initiate optimization from the input feature space.
-When the input and output dimensions differ, a downsampling operation is employed to ensure alignment.
-During inference, a scalar parameter $\alpha$ balances the contributions of the main branch and the input features prior to their combination.
+The architecture encompass a residual block featuring a zero-initialized, bias-free linear layer to commence optimization from the input feature space.
+When the input and output dimension differ, a downsampling operation is utilized to achieve alignment.
+During inference, a scalar parameter $\alpha$ balance the main branch and input features before combination.
 4.2 Isolating Content from Style with Augmented Prompts
 Despite progress in disentangling content and style via image augmentation, adequately altering all style factors in an image remains challenging due to the high dimensionality and complexity of style information in images.
-Achieving substantial style changes through augmentation—essential for complete disentanglement [41]—is difficult with existing image augmentation techniques.
+Achieving substantial style changes through augmentation, essential for complete disentanglement [41], is difficult with existing image augmentation techniques.
 On the contrary, text data inherently possesses high semanticity and logical structure, making it more amenable to property-wise manipulation compared to image data.
-To further explore the disentanglement of content, we propose Contrastive Learning with Augmented Prompts (CLAP).
+To further exploring the disentanglement of content, we propose Contrastive Learning with Augmented Prompts (CLAP).
 As depicted in Fig.
-2(b), CLAP employs an InfoNCE loss to train a disentangled network atop CLIP’s pre-trained text encoder, keeping the encoder’s gradients fixed—similar to Im.Aug.
-Leveraging the simpler structure of text, the template-based prompts previously utilized for synthesizing images now serve as the training text dataset, denoted by $\mathcal{D}_{t}$.
-Utilizing the same disentangled network as in Im.Aug, the learning objective of CLAP is defined as:
-$\mathbf{f}^{*}_{c}=\mathop{\mathrm{argmin}}_{\mathbf{f}_{c}}\mathop{\mathbb{E}}_{\{\mathbf{t}_{i}\}_{i=1}^{b}\in\mathcal{D}_{t}}\mathcal{L}(\mathbf{f}^{*}_{c}\circ\mathbf{f}^{*}_{t}(\mathbf{t});\{\mathbf{t}_{i},\mathbf{\tilde{t}}_{i}\}_{i=1}^{b},\tau)+\lambda\mathcal{L}(\mathbf{f}^{*}_{c}\circ\mathbf{f}^{*}_{t}(\mathbf{t});\{\mathbf{t}^{c}_{i},\mathbf{\tilde{t}}_{i}\}_{i=1}^{b},1),$ (4)
-where $\mathbf{f}^{*}_{t}$ denotes the pre-trained CLIP text encoder.
-The term $\mathbf{t}_{i}$ refers to a text prompt from $\mathcal{D}_{t}$, and $\mathbf{\tilde{t}}_{i}$ represents its augmented view, produced via prompt augmentation techniques.
-On the right-hand side of the equation, $\mathbf{t}_{i}^{c}$ denotes the class name associated with the text prompt $\mathbf{t}_{i}$.
-This strategy aims to enhance variation between prompt pairs, especially in cases where the text dataset $\mathcal{D}_{t}$ contains a limited number of samples.
-The hyperparameter $\lambda$ controls the relative weight of the second term in the overall loss function.
+2(b), CLAP employs an InfoNCE loss to train a disentangled network atop CLIP’s pre-trained text encoder, keeping the encoder’s gradients fixed, similar to Im.Aug.
+Leveraging the simpler structure of text, the template-based prompts previously utilized for synthesizing images now serve as the training text dataset, denoted by $\mathcal{D}_{\mathbf{t}}$.
+Utilizing the same disentangled network as in Im.Aug, the learning objective of CLAP is outlined as follows:
+$\mathbf{f^{*}_{c}}=\mathop{\mathrm{argmin}}_{\mathbf{f_{c}}}\mathop{\mathbb{E}}_{\{\mathbf{t}_{i}\}_{i=1}^{b}\in\mathcal{D}_{\mathbf{t}}}\mathcal{L}(\mathbf{f_{c}\circ f^{*}_{t}};\{\mathbf{t}_{i},\mathbf{\tilde{t}}_{i}\}_{i=1}^{b},\tau)+\lambda\mathcal{L}(\mathbf{f_{c}\circ f^{*}_{t}};\{\mathbf{t}^{c}_{i},\mathbf{\tilde{t}}_{i}\}_{i=1}^{b},1),$ (4)
+where $\mathbf{f}^{*}_{\mathbf{t}}$ denotes the pre-trained CLIP text encoder.
+The term $\mathbf{t}_{i}$ references a text prompt from $\mathcal{D}\mathbf{t}$, and $\mathbf{\tilde{t}}_{i}$ represents its augmented view, produced via prompt augmentation techniques.
+On the equation’s right side, $\mathbf{t}_{i}^{c}$ specifies the class name associated with the text prompt $\mathbf{t}_{i}$.
+This strategy aims to enhance variations between prompt pairs, especially in cases where the text dataset $\mathcal{D}_{\mathbf{t}}$ has a very limited number of samples.
+Here, $\lambda$ serves for adjusting the second term’s importance in the total loss function.
 All other symbols in Eq.
-4 retain their definitions as described earlier.
-After training, the learned disentangled network is seamlessly integrated with both of CLIP’s encoders to extract content representations, as shown in fig. 2(c).
-Prompt Augmentation
-To ensure text prompts undergo stylistic changes without compromising their content, we develop specific augmentation techniques for synthetic text prompts.
-Drawing inspiration from Easy Data Augmentation (EDA) [42], we adapt the Random Deletion (RD) and Random Swap (RS) techniques, customizing them for our structured prompt format.
-To avoid inadvertently altering the content—such as by introducing new object names or modifying the core semantics—we exclude random insertions and replacements.
-Our primary augmentation techniques include Object Size Deletion (OSD), Object Color Deletion (OCD), Image Type Deletion (ITD), Art Style Deletion (ASD), and Swapping Prompt Order (SPO).
-Each is applied with a fixed probability, as detailed in table 2.
-Additionally, for downstream datasets with few categories, we enrich the training set population using an additional augmentation method named IGN (Inserting Gaussian Noise).
-Following initialization protocols from prompt learning [48, 47], we insert zero-mean Gaussian noise with a standard deviation of 0.02 and noise length of 4 into the tokenized prompts.
+4 maintain their definitions as described earlier.
+After training, the learned disentangled network is seamlessly integrated with both of CLIP’s encoders to extract content representations, as depicted in Fig. 2(c).
+4.2.1 Prompt Augmentation
+To ensure text prompts undergo stylistic changes without compromising their content, we have developed specific augmentation techniques for synthetic text prompts.
+Drawing inspiration from Easy Data Augmentation (EDA) techniques [42], we adapted the Random Deletion (RD) and Random Swap (RS) techniques from EDA, customizing them to suit our prompt structure.
+To avoid inadvertently altering the content by introducing new object names or changing the core idea of a text prompt, our augmentation methods do not include random word insertions or replacements.
+Our primary augmentation techniques are Object Size Deletion (OSD), Object Color Deletion (OCD), Image Type Deletion (ITD), Art Style Deletion (ASD), and Swapping Prompt Order (SPO), each applied with a certain probability, as detailed in Tab.
+2.
+Additionally, for down-stream datasets with few categories, to rich the population of training samples, we use an additional augmentation, named IGN (Inserting Gaussian Noise).
+Following the initializing protocol of prompt learning methods [48, 47], we insert a zero-mean Gaussian noise with 0.02 standard deviation with a noise length equals to 4, to the tokenized prompts.
 Intuitively, these prompt augmentation methods parallel random masking techniques used in image augmentation [17, 6].
-However, prompt augmentations are often more effective and precise than their image-based counterparts.
-This is because prompt augmentations can directly target and remove specific style elements without affecting core content, whereas image masking operates at the pixel or patch level and may inadvertently distort content or fail to sufficiently alter style.
+However, prompt augmentations are more effective and precise than their image counterparts.
+This effectiveness arises because prompt augmentations can specifically target and eliminate a particular style element without impacting the content, whereas image masking, operating at the pixel or patch level, might inadvertently damage content information or lead to insufficient style changes.
 Table 2: Prompt augmentation techniques.
-Various augmented views are generated from an original text prompt using specific techniques: OSD (Object Size Deletion), OCD (Object Color Deletion), ITD (Image Type Deletion), ASD (Art Style Deletion), and SPO (Swapping Prompt Order).
-Prompt Variant OSD OCD ITD ASD SPO
-a realistic painting of a large red car a realistic painting of a red car a realistic painting of a large car a realistic of a large red car a painting of a large red car a large red car in a realistic painting
+Various augmented views are generated from an original text prompt using specific augmentation techniques: OSD (Object Size Deletion), OCD (Object Color Deletion), ITD (Image Type Deletion), ASD (Art Style Deletion), and SPO (Swapping Prompt Order).
+Original OSD OCD ITD ASD SPO
+a realistic a realistic a realistic a realistic a painting a large red
+painting of a  painting of a  painting of a of a of a car in a
+large red car red car large car  large red car  large red car  realistic painting
 5 Experiments
-We conduct three primary experiments to assess our method:
-  1. 1.
-    Zero-shot evaluation with diverse prompts to gauge performance and robustness under prompt perturbations;
-  2. 2.
-    Linear probe tests on few-shot samples to evaluate the efficacy of the learned representations in few-shot settings;
-  3. 3.
-    Adversarial attack assessments on zero-shot and one-shot classifiers to determine their resistance to adversarial threats.
-We further conduct an ablation study on hyperparameters, explore the impact of different prompt augmentation combinations and various sources of training prompts on CLAP’s performance, and replicate experiments across different CLIP model sizes.
+We conduct three primary experiments to assess our method: (1) zero-shot evaluation with diverse prompts to gauge zero-shot performance and its robustness to prompt perturbations; (2) linear probe tests on few-shot samples to evaluate the efficacy of the learned representations in few-shot settings; and (3) adversarial attack assessments on zero-shot and one-shot classifiers to determine their resistance to adversarial threats.
+We further conduct an ablative study on hyper-parameters, explore the impact of different prompt augmentation combinations and various sources of training prompts on CLAP’s performance, and replicate experiments across different CLIP model sizes.
 5.1 Experimental Setup
-Implementation
+Implementation.
 Im.Aug and CLAP are implemented using the ViT-B/16 CLIP model and executed on an NVIDIA RTX 3090 GPU.
 To ensure reproducibility, the random seed for all stochastic processes is fixed at 2024.
-Additional implementation details are provided in § A.1.
-Datasets
-CLAP is evaluated on four multi-domain datasets to assess its performance across diverse environments: PACS [23] (4 domains, 7 categories), VLCS [1] (4 domains, 5 categories), OfficeHome [37] (4 domains, 65 categories), and DomainNet [35] (6 domains, 345 categories).
-For conciseness, we present average results across domains for each dataset.
-Detailed results for individual domains are provided in § A.4.
-Compute Efficiency
+More information on implementation details is provided in Sec.
+0.A.1.
+Datasets.
+CLAP is assessed across four multi-domain datasets to examine its performance in varied environments: PACS [23] (4 domains, 7 categories), VLCS [1] (4 domains, 5 categories), OfficeHome [37] (4 domains, 65 categories), and DomainNet [35] (6 domains, 345 categories).
+For conciseness, we present average results across the domains for each dataset.
+Detailed experimental outcomes for each domain within these datasets are provided in Sec.
+0.A.4.
+Compute efficiency.
 CLAP demonstrates faster convergence and shorter training times compared to Im.Aug.
-Specifically, CLAP completes training on the PACS and VLCS datasets in approximately 11 minutes, OfficeHome in about 14 minutes, and DomainNet in roughly 47 minutes.
+For CLAP, training on the PACS and VLCS datasets is completed in roughly 11 minutes, OfficeHome in approximately 14 minutes, and DomainNet in about 47 minutes.
 In contrast, Im.Aug requires around 16 minutes for PACS and VLCS, 50 minutes for OfficeHome, and 3.3 hours for DomainNet.
-Both Im.Aug and CLAP retain CLIP’s inference efficiency due to the disentangled network’s compact two-layer MLP structure.
+Both Im.Aug and CLAP maintain CLIP’s inference efficiency due to the disentangled network’s efficient two-layer MLP structure.
 5.2 Main Results
-Zero-Shot Performance
-To assess zero-shot capabilities, CLAP is evaluated using three specific fixed prompts: ZS(C), which uses only the class name as "[class]"; ZS(PC), formatted as "a photo of a [class]"; and ZS(CP), structured as "a [class] in a photo".
-To further examine robustness, we introduce a dynamic prompt, ZS(NC), structured as "[noise][class]", where "[noise]" refers to Gaussian noise with mean 0 and standard deviation 0.02.
-As shown in table 3, CLAP outperforms both CLIP and Im.Aug across all prompts and datasets.
-Unlike the uniform improvement achieved by CLAP, Im.Aug yields inconsistent results.
-Notably, CLAP exhibits a substantial advantage over CLIP when evaluated with the dynamic ZS(NC) prompt, demonstrating its effectiveness in improving zero-shot performance through enhanced content disentanglement.
-Table 3: Zero-shot performance across prompt variants.
-Fixed prompts: "C" for [class], "CP" for a [class] in a photo, "PC" for a photo of a [class]; dynamic prompt "NC" is [noise][class] with Gaussian noise.
-CLAP consistently outperforms CLIP, while image augmentation (Im.Aug) shows mixed results.
-Prompt Method        PACS      VLCS OfficeHome DomainNet Overall
+5.2.1 Zero-Shot Performance
+To assess zero-shot capabilities, CLAP undergoes evaluation using three specific fixed prompts: ZS(C), utilizing only the class name within "[class]"; ZS(PC), with the format "a photo of a [class]"; and ZS(CP), structured as "a [class] in a photo".
+To thoroughly examine zero-shot proficiency, a dynamic prompt, ZS(NC), formatted as "[noise][class]", is also used, where "[noise]" signifies the introduction of Gaussian noise characterized by a mean of 0 and a standard deviation of 0.02.
+As presented in Tab.
+3, CLAP surpasses both CLIP and Im.Aug across all evaluated prompts for every dataset.
+Unlike the uniform enhancement in zero-shot performance CLAP achieves over CLIP, Im.Aug displays inconsistent results.
+A closer examination reveals CLAP’s superiority over CLIP is especially significant for the dynamic ZS(NC) prompt.
+This demonstrates CLAP’s effectiveness in significantly improving zero-shot performance compared to the original CLIP representations.
+Table 3: Zero-shot results across three distinct prompts: "C" for "[class]", "CP" for "a [class] in a photo", "PC" for "a photo of a [class]", and a dynamic prompt "NC" for "[noise][class]" showcase that CLAP consistently outperforms CLIP’s zero-shot performance across all datasets, whereas image augmentation exhibits mixed outcomes.
+Prompt Method Zero-shot performance, avg. top-1 acc. (%) (↑)
+ PACS  VLCS Off.Home  Dom.Net Overall
 ZS(C) CLIP 95.7 76.4 79.8 57.8 77.4
@@
 CLAP 97.2 81.0 73.5 52.6 76.1
-To evaluate robustness to prompt perturbations, we analyze the variance in zero-shot performance across prompts using the range ($R$) and standard deviation ($\delta$) from ZS(C), ZS(CP), and ZS(PC).
-We also measure the performance drop ($\Delta_{(NC)}$) from ZS(C) to ZS(NC) to assess resilience to noisy prompts.
-As shown in table 4, CLAP significantly reduces zero-shot variance, exhibiting lower $\delta$ and $R$ values, and a smaller drop in performance under noise, compared to both Im.Aug and CLIP.
-While Im.Aug offers modest variance reduction, it is clearly less effective than CLAP.
-These results demonstrate CLAP’s superior robustness in maintaining consistent performance across prompt variations.
-Table 4: Prompt robustness analysis.
-CLAP more effectively reduces zero-shot performance variance across prompts than image augmentation.
-$R$ and $\delta$ denote the range and standard deviation across ZS(C), ZS(CP), and ZS(PC); $\Delta_{(NC)}$ indicates the performance drop from ZS(C) to ZS(NC).
-Lower values imply greater robustness.
-Metric Method        PACS      VLCS OfficeHome DomainNet Overall
+In assessing the model’s robustness to prompt perturbations, we examine the variances in zero-shot performance across different prompts by analyzing the range ($R$) and standard deviation ($\delta$) of results derived from ZS(C), ZS(CP), and ZS(PC).
+Additionally, we investigate the decline ($\Delta_{(NC)}$) in performance from ZS(C) to ZS(NC) as a broad indicator of resilience to noised prompts.
+As presented in Tab.
+4, CLAP significantly reduces the variance in zero-shot performance across various testing prompts, evidenced by markedly lower values of $\delta$ and $R$, and a less pronounced decrease in performance with a noised prompt, in contrast to Im.Aug and the baseline representations of CLIP.
+Although Im.Aug aids in reducing performance variance to some extent, its efficacy is notably inferior to that of CLAP.
+These findings highlight CLAP’s enhanced robustness in maintaining consistent zero-shot performance across a diverse array of prompts.
+Table 4: CLAP more effectively reduces zero-shot performance variance across prompts than image augmentation, with $R$ and $\delta$ indicating the range and standard deviation for ZS(C), ZS(CP), and ZS(PC).
+The decrease $\Delta_{(NC)}$ from ZS(C) to ZS(NC) highlights CLAP’s enhanced robustness against prompt perturbations.
+Metric Method Performance variance, avg. top-1 acc. (%) (↓)
+ PACS  VLCS Off.Home  Dom.Net Overall
 $R$ CLIP 0.9 6.1 3.1 0.8 2.7
@@
 CLAP 0.0 1.6 7.5 6.1 3.8
-Few-Shot Performance
-We conduct evaluations using 1-shot, 4-shot, 8-shot, and 16-shot linear probes across each domain within the four datasets.
+5.2.2 Few-Shot Performance
+We conduct evaluations of 1-shot, 4-shot, 8-shot, and 16-shot linear probes across each domain within the four datasets.
 As illustrated in Fig.
 4, CLAP significantly outperforms both CLIP and Im.Aug in few-shot learning scenarios.
-Notably, in the 1-shot setting, CLAP achieves performance improvements over the linear-probe CLIP model by margins of +10%, +3.5%, +2.5%, and +1.5% on the PACS, VLCS, OfficeHome, and DomainNet datasets, respectively.
-These improvements are especially pronounced when compared to the gains observed with Im.Aug, highlighting CLAP’s superior effectiveness in few-shot learning.
-Detailed quantitative results are provided in § A.4.
+Notably, in the 1-shot setting CLAP achieves performance improvements over the linear-probe CLIP model by margins of +10%, +3.5%, +2.5%, and +1.5% on the PACS, VLCS, OfficeHome, and DomainNet datasets, respectively.
+These improvements are especially significant in comparison to the gains observed with Im.Aug counterparts, underpinning CLAP’s efficacy in few-shot scenarios.
+For detailed quantitative results, please refer to Sec.
+0.A.4.2.
 Figure 4: Few-shot linear probe comparisons of image-encoder features show that CLAP enhances CLIP’s few-shot performance more effectively than Im.Aug.
-"ZS" indicates zero-shot performance using a [class] prompt.
-Adversarial Performance
-To evaluate adversarial robustness, we test both zero-shot (ZS(C)) and one-shot classifiers against three popular adversarial attack methods: FGSM [15], PGD [30], and CW [5].
-Adversarial samples are generated with 1 iteration for FGSM and 20 iterations for both PGD and CW, using an $\epsilon$ of 0.031.
-As shown in table 5, classifiers leveraging CLAP representations exhibit significantly greater robustness than those based on CLIP.
-Across the four datasets, CLAP’s zero-shot and one-shot classifiers outperform CLIP by +7.6% and +8.5% against FGSM, +1.0% and +11.7% against PGD-20, and +1.1% and +2.3% against CW-20, respectively.
-These gains surpass the improvements achieved by Im.Aug: +4.4% and +4.6% (FGSM), +0.3% and +6.2% (PGD-20), and 0% and +1.3% (CW-20).
-These results indicate that CLAP enhances robustness against adversarial attacks in both zero-shot and one-shot scenarios.
-Table 5: Adversarial robustness evaluation.
-Both Im.Aug and CLAP improve the robustness of CLIP’s zero-shot ([class]) and 1-shot classifiers under adversarial attacks (FGSM, PGD-20, CW-20).
-CLAP consistently achieves the highest resilience across all datasets.
-Setting Method FGSM PGD-20 CW-20 Avg.
+In the accompanying figure, "ZS" indicates the zero-shot performance using a "[class]" prompt.
+5.2.3 Adversarial Performance
+To assess adversarial robustness, zero-shot (ZS(C)) and one-shot classifiers are evaluated against prominent adversarial attack methods, such as FGSM [15], PGD [30], and CW [5], by generating adversarial samples for testing.
+For FGSM, 1 adversarial iteration is employed, whereas for PGD and CW, 20 iterations are used, all with an epsilon of 0.031.
+As indicated in Tab.
+5, classifiers utilizing CLAP representations demonstrate superior resilience to these adversarial attacks compared to those based on CLIP representations.
+Across the four datasets, CLAP’s zero-shot and 1-shot classifiers surpass CLIP by margins of +7.6% and +8.5% against FGSM, +1.0% and +11.7% against PGD-20, and +1.1% and +2.3% against CW-20, respectively.
+These figures notably exceed the performance improvements of +4.4% and +4.6% against FGSM, +0.3% and +6.2% against PGD-20, and 0% and +1.3% against CW-20 achieved by Im.Aug.
+The result suggests that CLAP efficiently enhances robustness against adversarial attacks in both zero-shot and one-shot scenarios.
+Table 5: Image augmentation and CLAP both enhance CLIP’s zero-shot with the "[class]" prompt and 1-shot robustness against adversarial attacks, with CLAP showing greater improvements.
+Setting Method Avg. top-1 acc. (%) under adversarial attacks(↑)
+FGSM PGD-20 CW-20  Avg.
  PACS  VLCS  O.H.  D.N.  PACS  VLCS  O.H.  D.N.  PACS  VLCS  O.H.  D.N.
@@
 5.3 More Analysis
-t-SNE Visualization
+5.3.1 t-SNE Visualization
 In our t-SNE visualizations, we examine the representations of CLIP, Im.Aug, and CLAP for all images within the Art Painting domain of the PACS dataset.
@@
 Figure 5: t-SNE visualizations of all images in the Art Painting of PACS dataset show CLAP outperforms the original CLIP and Im.Aug, with clearer inter-class distinctions and tighter intra-class clusters.
-Ablations
+5.3.2 Ablations
 In Fig.
@@
 CLAP continuously enhance CLIP’s performance throughout the tested values.
-Prompt Augmentation Combinations
-We explore diverse combinations of our tailored prompt augmentation techniques and examine the impact of Easy Data Augmentation (EDA) methods [42] on the VLCS dataset.
-All evaluated techniques demonstrate that CLAP provides consistent improvements over CLIP.
-Further details are presented in § A.2.
-Prompt Sources
-We assess the influence of different training prompt sources on performance, using the VLCS dataset and incorporating EDA techniques.
-The evaluated prompt formats include: (1) our template-based prompts; (2) LLM-generated prompts from ChatGPT-3.5 [3], with generation details in § B.2; (3) prompts of the form "a [random] style of [class]," where "[random]" is sampled from a random word generator [^1] https://github.com/vaibhavsingh97/random-word; and (4) prompts generated by the PromptStyler method [9].
-The results suggest that training with simpler, more structured prompts tends to yield better performance.
-Detailed quantitative results are reported in § A.3.
-Experiments on Different Model Scales
-We repeat our zero-shot evaluations using larger pre-trained CLIP models, specifically ViT-L/14 and ResNet50x16.
-Across both model scales, CLAP consistently improves zero-shot performance while also reducing variance.
-These findings reaffirm CLAP’s scalability and robustness in enhancing CLIP representations.
-Quantitative results are provided in §§ B.3 and B.4.
+5.3.3 Prompt Augmentation Combinations
+We explore diverse combinations of our tailored prompt augmentation methods and examine Easy Data Augmentation (EDA) techniques [42] on the VLCS dataset.
+Each tested technique showcases CLAP’s enhancements over CLIP, with details available in Sec.
+0.A.2.
+5.3.4 Prompt Sources
+We assess the impact of different training prompt formats, originating from various synthetic sources, on the performance of the VLCS dataset, incorporating EDA techniques.
+Our evaluation includes our template-based prompts, LLM-generated prompts by ChatGPT-3.5 [3] (with the generation process detailed in Sec.
+0.B.2 ), prompts structured as "a [random] style of [class]," where "[random]" is filled with terms from a random word generator [^2] https://github.com/vaibhavsingh97/random-word, and prompts produced using the PromptStyler method [9].
+The findings indicate that the training prompts with simpler forms tend to yield better performance, with detailed quantitative results presented in Sec.
+0.A.3.
+5.3.5 Experiments on Different Model Scales
+In our repeated experiments assessing zero-shot performance on the ViT-L/14 and ResNet50x16 pre-trained with CLIP, we consistently find that CLAP improves zero-shot performance while also reducing performance variances.
+This consistent observation underscores CLAP’s effectiveness in enhancing the quality of CLIP representations.
+For quantitative details supporting these findings, please see the Appendix 0.C.
 6 Conclusion
-To enhance the representations of pre-trained CLIP-like models, this study investigates the disentanglement of latent content variables.
-Adopting a causal perspective on the generative processes underlying vision-language data, we demonstrate that training a disentangled network in one modality enables effective disentanglement across both modalities.
-Leveraging the semantic richness and logical structure of text data, we find that disentanglement is more readily achieved in the language modality through targeted text augmentation.
-Based on these insights, we propose CLAP (Contrastive Learning with Augmented Prompts), a method designed to acquire disentangled vision-language content features.
-Extensive experiments validate the effectiveness of CLAP, showing substantial improvements in zero-shot and few-shot classification tasks, along with enhanced robustness to input perturbations.
-We hope this work will motivate further research into disentangled representation learning within multimodal frameworks.
-Appendix A More on Experiments with ViT-B/16
-A.1 Implementation Details
-This section provides detailed implementation settings for our experiments using the CLIP pre-trained ViT-B/16 model.
+To enhance pre-trained CLIP-like models, this study delves into disentangling latent content variables.
+Through a causal analysis of the underlying generative processes of vision-language data, we discover that training a disentangled network in one modality can effectively disentangle content across both modalities.
+Given the high semantic nature of text data, we identify that disentanglement is more achievable within the language modality through text augmentation interventions.
+Building on these insights, we introduce CLAP (Contrastive Learning with Augmented Prompts) to acquire disentangled vision-language content features.
+Comprehensive experiments validate CLAP’s effectiveness, demonstrating significant improvements in zero-shot and few-shot performance, and enhancing robustness against perturbations.
+We anticipate that our work will inspire further exploration into disentangling latent variables within vision-language models.
+Appendix 0.A More on Experiments with ViT-B/16
+0.A.1 Implementation Details
+In this section, we detail the implementation of our experiments utilizing the CLIP pre-trained ViT-B/16 model:
 Network.
-The network’s output dimension is set to match the 512-dimensional CLIP feature space, thereby eliminating the need for feature downsampling.
-The latent dimension is tailored for each dataset: 256 for PACS, 448 for OfficeHome, and 512 for VLCS and DomainNet, to accommodate differences in category counts and dataset complexity.
-The weight parameter $\alpha$ is adjusted per dataset as follows: 0.208 for PACS, 0.056 for VLCS, 0.14 for OfficeHome, and 0.2 for DomainNet.
-During training, $\alpha$ is consistently set to 1.
+The network’s output dimension is aligned with the 512-dimensional CLIP features, thereby obviating the need for input feature downsampling.
+The latent dimensions are tailored to each dataset: 256 for PACS, 448 for OfficeHome, and 512 for VLCS and DomainNet, to accommodate the variety of categories and complexity of datasets.
+The weight parameter $\alpha$ is adjusted to 0.208 for PACS, 0.056 for VLCS, 0.14 for OfficeHome, and 0.2 for DomainNet, while it is consistently maintained at 1 throughout the training phase.
 Training CLAP.
-We use consistent training parameters across all datasets.
-The Adam optimizer is employed with a learning rate of 0.0001.
-Training is limited to 8,000 steps, with the average loss evaluated every 480 steps, and early stopping applied if the loss does not decrease by at least 0.01 over five checkpoints.
-Batch sizes are set to 8 for PACS and VLCS, 96 for OfficeHome, and 384 for DomainNet.
-The temperature parameter $\tau$ is configured to 0.5 for PACS and VLCS, and 0.3 for OfficeHome and DomainNet.
-The loss coefficient $\lambda$ is set to 1 for PACS and VLCS, and 0.0001 for OfficeHome and DomainNet, reflecting the fewer number of classes in the former two datasets.
-Prompt augmentations OSD, OCD, and SPO are applied across all datasets with a probability of 0.5.
-For PACS and VLCS, Gaussian noise (mean 0, standard deviation 0.02) is randomly inserted at the beginning, middle, or end of the augmented prompts to diversify the training samples.
-For linear probe evaluations in the few-shot setting, we use L2 normalization and cross-entropy loss, training over 1,000 epochs with a batch size of 32.
-Early stopping is used with a patience of 10 epochs and a loss improvement threshold of 0.001.
+Training parameters are consistent across datasets, employing the Adam optimizer with a learning rate of 0.0001, limiting training to 8,000 steps with checking the average loss every 480 steps, and instituting early stopping after five checkpoints without a loss decrease of at least 0.01.
+Batch sizes are adjusted to 8 for PACS and VLCS, 96 for OfficeHome, and 384 for DomainNet, with the temperature parameter $\tau$ set at 0.5 for PACS and VLCS, and 0.3 for OfficeHome and DomainNet.
+The loss coefficient $\lambda$ is set to 1 for PACS and VLCS, and 0.0001 for OfficeHome and DomainNet, due to the first two datasets have less classes.
+Prompt augmentations, OSD+OCD+SPO, are applied across datasets all with a 0.5 probability.
+For the PACS and VLCS datasets, Gaussian noise with a zero mean and a standard deviation of 0.02 is randomly inserted at the beginning, middle, or end of the augmented-view prompts to enrich the training samples.
+In the linear probe evaluations for few-shot analysis, L2 normalization and cross-entropy loss are utilized for training over 1,000 epochs with a batch size of 32, incorporating early stopping with a patience threshold of 10 epochs and a loss decrease criterion of 0.001.
 Training Im.Aug.
-We train a disentangled network using image augmentations and the InfoNCE loss with a temperature parameter $\tau$ set to 0.5.
-Image augmentations include random cropping (scale $\in[0.64,1.0]$) and color distortion (brightness = 0.5, hue = 0.3), each applied with a probability of 0.5.
-All other training and inference configurations for Im.Aug are consistent with those used for CLAP across the datasets.
-A.2 Prompt Augmentation Combinations
-In table 6, we evaluate various combinations of our tailored prompt augmentation techniques alongside Easy Data Augmentation (EDA) [42] on the VLCS dataset.
-Across all configurations, CLAP consistently enhances CLIP’s performance while effectively reducing performance variance across different prompt conditions.
-Among the tested combinations, OSD+OCD+SPO+IGN (①②⑤⑥) yields the highest average zero-shot accuracy and the lowest variance, outperforming all other augmentation strategies, including EDA.
-Notably, even without the inclusion of random noise (IGN), CLAP achieves substantial improvements over CLIP, particularly in mitigating performance degradation under noisy prompt conditions, as reflected by a significantly lower $\Delta_{(NC)}$.
-Table 6: Effect of prompt augmentation combinations on VLCS.
-We assess the impact of various augmentations: OSD (①), OCD (②), ITD (③), ASD (④), SPO (⑤), and IGN (⑥).
-ZS(Avg.) denotes the average zero-shot accuracy across four different inference prompts.
-CLAP improves both accuracy and robustness, with the best results achieved using the combination ①②⑤⑥.
-Metric CLIP (base) EDA ①②③④⑤⑥ ①②③④⑤ ①②③④⑥ ①②③④ ③④⑤ ①②⑤⑥
+We train a disentangled network using image augmentation, applying the InfoNCE loss with a temperature parameter $\tau$ set to 0.5.
+This include image augmentation techniques, image cropping ($scale\in[0.64,1.0]$) and color distortion ($brightness=0.5,hue=0.3$), each with a probability of 0.5.
+Other training and inference configurations for Im.Aug are consistent with those used for CLAP across all datasets.
+0.A.2 Prompt Augmentation Combinations
+In Tab.
+6, we explore different combinations of our tailored prompt augmentation techniques and EDA (Easy Data Augmentation) [42] techniques on the VLCS dataset.
+Each combination demonstrates CLAP’s effectiveness in enhancing CLIP’s performance and reducing performance disparities.
+The combination of OSD+OCS+SPO+IGN achieves the highest average accuracy and the least variance, outperforming the EDA techniques.
+Notably, even without incorporating random noise in the augmentations, CLAP significantly surpasses CLIP in handling perturbations on prompts, as evidenced by the largely reduced $\Delta_{(NC)}$.
+Table 6: We evaluate prompt augmentation combinations on the VLCS dataset: OSD (①), OCD (②), ITD (③), ASD (④), SPO (⑤), and IGN (⑥).
+ZS(Avg.) shows average zero-shot accuracy acoss four distinct inference prompts.
+CLAP boosts CLIP’s accuracy and reduces variances, with ①②⑤⑥ as the optimal combination.
+Metrics CLIP (base) Avg. top-1 acc. (%) of different augmentations
+EDA ①②③ ①②③ ①②③ ①②③ ③④⑤ ①②⑤
+④⑤⑥ ④⑤ ④⑥
 ZS(Avg.) (↑) 77.3 81.6 82.0 80.1 82.0 79.6 82.1 82.6
@@
 $\Delta_{(NC)}$ (↓) 8.1 2.3 1.7 3.0 1.8 3.4 2.0 1.6
-A.3 Prompt Sources
-In table 7, we evaluate the impact of different training prompt formats, sourced from various synthetic origins, on the performance of CLAP using the VLCS dataset.
-All experiments in this subsection are conducted with EDA (Easy Data Augmentation) techniques applied during training.
-The evaluated prompt formats include: Template, representing our structured template-based prompts; LLM, referring to prompts generated using ChatGPT-3.5 [3] (see § B.2 for generation details); Random, which uses the format "a [random] style of [class]", where "[random]" is filled using a random word generator; and Prm.Stl., which denotes vectorized prompts produced by the PromptStyler method [9].
-Table 7: Effect of training prompt sources on CLAP performance (VLCS).
-We train CLAP using EDA-augmented prompts from various sources.
-All prompt types improve CLIP’s baseline zero-shot performance.
-Simpler prompt formats—Random and Template—achieve the highest accuracy and robustness.
-Metric CLIP (base) LLM Random Prm.Stl. Template
+0.A.3 Prompt Sources
+In Tab. 7, we examine the effects of various training prompt formats, sourced from different synthetic origins, on the VLCS dataset performance, utilizing
+Table 7: We employ EDA augmentation to train CLAP with diverse prompt sources on the VLCS dataset.
+Each prompt source contributes to improvements in CLIP’s zero-shot performance, with "Random" and "Template" prompts, in their simpler forms, yielding better outcomes.
+Metrics CLIP (base) Avg. top-1 acc. (%) of different sources
+  LLM  Random Prm.Stl. Template
 ZS(Avg.) (↑) 77.3 78.2 81.6 81.2 81.6
@@
 $\Delta_{(NC)}$ (↓) 8.1 3.3 2.3 3.0 2.3
-The results indicate that CLAP trained with any of the evaluated prompt sources leads to consistent improvements over the original CLIP model.
-Interestingly, despite the sophistication of LLM-generated and PromptStyler-based prompts, the simpler Random and Template prompts outperform them in both accuracy and robustness.
-It is worth noting, however, that while these alternative prompt formats paired with EDA enhance zero-shot performance, they do not surpass the performance achieved by prompt augmentations specifically optimized for the template-based format (as reported in § A.2).
-A.4 Detailed Results on ViT-B/16
-Details on Zero-Shot Evaluations
-We present domain-level zero-shot performance using various prompts across each dataset in table 8.
-CLAP consistently improves upon CLIP across all prompt variants and domains.
-Notably, as CLAP exclusively uses text data for training, it retains CLIP’s domain generalization capabilities—an outcome of CLIP’s large-scale training corpus.
-Rather than compromising this generalization, CLAP enhances it through more effective disentanglement of content information, leading to consistent zero-shot gains across all evaluated domains.
-Table 8: Domain-level zero-shot results with ViT-B/16.
-We report top-1 accuracy (%) using four different prompts: ZS(C), ZS(CP), ZS(PC), and ZS(NC), across dataset domains.
-CLAP improves zero-shot performance in all settings.
-Datasets Domain-level avg. top-1 acc. (%) of zero-shot performance usig ViT-B/16 (↑)
+EDA techniques.
+The prompt formats are defined as follows: "Template" refers to the template-based prompts fundamental to our primary approach; "LLM" designates prompts created by ChatGPT-3.5 [3], with the generation process elaborated in Sec.
+0.B.2; "Random" describes prompts formatted as "a [random] style of [class]," with "[random]" being replaced by terms from a random word generator; and "Prm.Stl." indicates vectorized prompts generated through PromptStyler[9].
+Our experimental results indicate that CLAP, when trained across these varied prompt formats, enhances the performance of CLIP.
+Notably, despite the complex generation mechanisms of "LLM" and "Prm.Stl." prompts, the simpler, random-styled and template-based prompts demonstrate superior efficacy.
+However, it is important to highlight that the improvements attributed to these diverse prompt formats, trained with EDA, do not surpass the best performance of the prompt augmentations tailored for template-based prompts.
+0.A.4 Detailed Results on ViT-B/16
+0.A.4.1 Details on Zero-Shot Evaluations
+We present the domain-level zero-shot performance with various prompts across each dataset in Tab.
+8.
+CLAP consistently enhances CLIP’s zero-shot performance across these different prompts.
+Given that CLAP exclusively utilizes text data for training, it does not compromise CLIP’s inherent ability to generalize across domains, which is acquired from its extensive training dataset.
+Rather, by achieving a more effective disentanglement of content, it unequivocally enhances CLIP’s zero-shot performance across all dataset domains.
+Table 8: Domain-level zero-shot results of the ViT-B/16 model on the test datasets.
+Dataset Domains Domain-level avg. top-1 acc. (%) of zero-shot performance usig ViT-B/16 (↑)
 ZS(C) ZS(CP) ZS(PC) ZS(NC)
@@
 V. 73.9 77.1 84.9 84.8 85.4 86.0 86.1 85.6 86.2 68.9 70.3 82.9
-O.H. A. 80.5 79.0 81.8 80.1 76.0 81.6 83.2 78.7 83.2 73.0 69.2 73.6
+OfficeHome A. 80.5 79.0 81.8 80.1 76.0 81.6 83.2 78.7 83.2 73.0 69.2 73.6
 C. 64.6 59.6 66.4 63.7 58.9 65.4 68.1 61.9 69.0 57.0 52.0 60.4
@@
 R. 88.0 85.9 88.5 87.6 84.8 87.7 89.8 87.2 90.0 79.0 76.5 81.1
-D.N. C. 71.0 64.3 71.9 70.5 62.1 72.0 71.3 63.4 72.8 63.2 53.9 64.6
+DomainNet C. 71.0 64.3 71.9 70.5 62.1 72.0 71.3 63.4 72.8 63.2 53.9 64.6
 I. 48.6 40.5 50.6 47.7 40.7 49.5 47.8 40.0 50.5 42.9 35.0 45.1
@@
 S. 63.1 56.1 63.7 62.2 55.0 63.1 63.4 56.4 64.4 55.7 47.5 57.6
-Table 9: Domain-level few-shot results with ViT-B/16.
-We report top-1 accuracy (%) for 1-shot through 32-shot settings across dataset domains.
-CLAP outperforms both CLIP and Im.Aug, particularly in low-shot scenarios.
-Datasets Domain-level avg. top-1 acc. (%) of few-shot performance of ViT-B/16 (↑)
+Table 9: Domain-level few-shot results of the ViT-B/16 model using the test datasets.
+Dataset Domains Domain-level avg. top-1 acc. (%) of few-shot performance of ViT-B/16 (↑)
 1-shot 4-shot 8-shot 16-shot 32-shot
@@
 V. 50.9 53.4 59.0 64.5 66.7 76.1 75.4 74.1 78.7 72.6 73.9 77.7 85.7 86.1 87.9
-O.H. A. 42.6 45.1 43.9 76.8 77.6 77.7 84.8 86.0 85.5 91.8 92.1 92.1 97.4 97.5 97.5
+OfficeHome A. 42.6 45.1 43.9 76.8 77.6 77.7 84.8 86.0 85.5 91.8 92.1 92.1 97.4 97.5 97.5
 C. 40.1 45.0 43.8 69.9 70.2 70.5 75.8 75.9 76.6 81.6 81.6 81.6 89.0 89.0 89.2
@@
 R. 58.4 59.3 59.4 81.7 83.1 82.9 89.7 89.5 89.9 92.9 92.7 93.2 95.8 95.8 95.8
-D.N. C. 42.1 43.6 43.8 66.8 67.5 67.8 74.2 74.3 74.6 78.5 78.6 78.8 82.8 82.8 82.7
+DomainNet C. 42.1 43.6 43.8 66.8 67.5 67.8 74.2 74.3 74.6 78.5 78.6 78.8 82.8 82.8 82.7
 I. 19.5 20.8 21.0 38.5 39.3 39.7 46.7 47.0 47.3 53.2 53.2 53.6 60.0 59.9 60.1
@@
 S. 33.1 33.9 34.8 56.2 56.6 57.2 62.9 62.9 63.7 67.8 67.7 68.1 72.5 72.3 72.6
-Details on Few-Shot Evaluations
-We present detailed quantitative results of few-shot performance in table 9.
-CLAP consistently enhances few-shot capabilities, demonstrating improvements across all evaluated domains and test datasets.
-Details on Adversarial Evaluations
-In table 10, we report domain-level adversarial performance for PACS, VLCS, OfficeHome, and DomainNet.
-CLAP improves both zero-shot and one-shot robustness across all domains.
-Although Im.Aug strengthens one-shot robustness, its impact on zero-shot adversarial performance remains inconsistent.
-Details on Ablative Analysis
-table 11 presents the detailed results of zero-shot performance using various combinations of prompt augmentation strategies.
-Additionally, table 12 summarizes the effects of key hyperparameters—including temperature $\tau$, latent dimension size, and weight coefficient $\alpha$—evaluated across individual domains.
-These findings confirm that CLAP maintains effectiveness across a broad range of hyperparameter configurations.
-Table 10: Domain-level adversarial results (ViT-B/16).
-Average top-1 accuracy (%) under adversarial attacks: FGSM, PGD-20, and CW-20, evaluated in both zero-shot (ZS-C) and one-shot settings across domains of PACS, VLCS, OfficeHome, and DomainNet.
-Datasets FGSM PGD-20 CW-20
+0.A.4.2 Details on Few-Shot Evaluations
+We display the quantitative results of few-shot performance in Tab. 9. CLAP consistently enhances the few-shot capabilities, showcasing improvements across test datasets at a closer domain level.
+0.A.4.3 Details on Adversarial Evaluations
+In Tab.
+10, we detail our adversarial performance evaluations for PACS, VLCS, OfficeHome, and DomainNet, respectively.
+CLAP enhances both zero-shot and one-shot performance across all domains of the tested datasets.
+While Im.Aug boosts one-shot robustness against adversarial tasks, its impact on zero-shot adversarial robustness is inconsistent.
+0.A.4.4 Details on Ablative Analysis
+In Tab.
+11, we provide detailed results from our analysis on zero-shot performance using various combinations of prompt augmentations.
+Additionally, in Tab.
+12, we present the outcomes of our ablative studies focusing on the hyperparameters $\tau$, latent dimension, and $\alpha$, respectively, each evaluated domain-wise.
+The results indicate that CLAP is effective across a wide range of hyperparameters.
+Table 10: Domain-level results under adversarial attacks of ViT-B/16 on the datasets.
+Dataset Domains Domain-level avg. top-1 acc. (%) under adversarial attackings using ViT-B/16 (↑)
+FGSM PGD-20 CW-20
 ZS-C 1-shot ZS-C 1-shot ZS-C 1-shot
 CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP CLIP Im.Aug CLAP
-PACS A 76.3 79.3 79.3 61.2 78.0 87.3 1.7 2.2 1.8 16.0 42.1 63.1 1.5 2.0 2.3 0.5 1.1 1.7
-C 94.9 95.0 94.0 66.5 84.2 95.1 33.3 37.7 35.6 33.3 57.2 86.1 28.8 34.0 33.2 11.9 23.6 31.8
-P 91.6 90.3 91.7 67.4 80.8 92.1 5.7 7.0 6.7 27.1 55.0 69.8 4.7 4.9 5.8 0.7 2.7 4.1
-S 84.5 87.5 89.8 71.6 74.6 83.8 75.8 78.4 79.2 63.0 66.3 74.6 74.5 76.8 77.9 62.7 65.4 70.3
-VLCS C 55.3 53.8 55.5 25.8 28.8 25.3 4.4 5.1 4.7 2.0 5.2 2.5 2.9 3.1 3.5 0.7 1.2 1.0
-L 49.4 45.5 50.6 27.0 32.6 30.4 15.2 14.9 16.0 6.4 8.9 8.0 12.4 11.2 13.0 6.1 8.3 7.7
-S 61.7 58.1 62.5 48.0 46.9 51.6 13.2 13.9 14.0 8.6 10.7 10.0 9.2 8.8 10.2 8.3 7.9 8.4
-V 65.3 63.2 65.6 36.5 40.1 41.0 7.5 7.9 7.9 5.3 9.4 8.9 5.2 4.8 5.6 2.9 2.8 2.9
-O.H. A 55.3 53.8 55.5 25.8 28.8 25.3 4.4 5.1 4.7 2.0 5.2 2.5 2.9 3.1 3.5 0.7 1.2 1.0
-C 49.4 45.5 50.6 27.0 32.6 30.4 15.2 14.9 16.0 6.4 8.9 8.0 12.4 11.2 13.0 6.1 8.3 7.7
-P 61.7 58.1 62.5 48.0 46.9 51.6 13.2 13.9 14.0 8.6 10.7 10.0 9.2 8.8 10.2 8.3 7.9 8.4
-R 65.3 63.2 65.6 36.5 40.1 41.0 7.5 7.9 7.9 5.3 9.4 8.9 5.2 4.8 5.6 2.9 2.8 2.9
-D.N. C 57.8 50.9 58.8 33.3 34.3 35.0 21.6 18.7 22.8 18.4 19.6 20.0 15.8 12.5 16.6 7.0 7.5 7.8
-I 35.8 28.0 37.0 12.2 13.3 13.2 6.1 3.7 6.7 4.6 5.3 5.1 3.3 1.9 3.7 0.9 0.9 0.9
-P 43.9 39.0 44.3 18.4 20.6 20.3 3.1 2.8 3.3 8.6 10.4 9.9 1.8 1.3 1.9 0.3 0.3 0.3
-Q 12.9 10.3 13.2 10.9 10.8 11.1 8.4 6.8 8.6 5.4 5.4 5.6 7.1 5.4 7.4 4.9 4.8 5.1
-R 62.1 55.9 62.4 34.5 35.9 36.5 7.1 6.5 7.5 17.6 19.7 19.6 4.5 3.4 4.7 1.2 1.4 1.4
-S 49.1 43.3 49.7 25.7 26.0 27.5 17.8 15.5 18.6 13.6 14.4 15.1 13.4 10.2 13.9 5.0 5.2 5.6
-Appendix B Data Synthesis
-B.1 Synthetic Image Generation
-We utilize the Stable Diffusion v2.1 model [39] for generating synthetic image data used in our comparative experiments.
-Specifically, we adopt the implementation provided via the Stable Diffusion v2-1 Model Card on Hugging Face [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1.
-For each class across the four benchmark datasets, a total of 480 images are synthesized by applying our structured template prompts as input to the diffusion model.
-All generated images are rendered at a resolution of $512\times 512$ pixels.
-Representative examples of the synthetic images, along with their corresponding text prompts, are presented in fig.
+PACS A. 76.3 79.3 79.3 61.2 78.0 87.3 1.7 2.2 1.8 16.0 42.1 63.1 1.5 2.0 2.3 0.5 1.1 1.7
+C. 94.9 95.0 94.0 66.5 84.2 95.1 33.3 37.7 35.6 33.3 57.2 86.1 28.8 34.0 33.2 11.9 23.6 31.8
+P. 91.6 90.3 91.7 67.4 80.8 92.1 5.7 7.0 6.7 27.1 55.0 69.8 4.7 4.9 5.8 0.7 2.7 4.1
+S. 84.5 87.5 89.8 71.6 74.6 83.8 75.8 78.4 79.2 63.0 66.3 74.6 74.5 76.8 77.9 62.7 65.4 70.3
+VLCS C. 55.3 53.8 55.5 25.8 28.8 25.3 4.4 5.1 4.7 2.0 5.2 2.5 2.9 3.1 3.5 0.7 1.2 1.0
+L. 49.4 45.5 50.6 27.0 32.6 30.4 15.2 14.9 16.0 6.4 8.9 8.0 12.4 11.2 13.0 6.1 8.3 7.7
+S. 61.7 58.1 62.5 48.0 46.9 51.6 13.2 13.9 14.0 8.6 10.7 10.0 9.2 8.8 10.2 8.3 7.9 8.4
+V. 65.3 63.2 65.6 36.5 40.1 41.0 7.5 7.9 7.9 5.3 9.4 8.9 5.2 4.8 5.6 2.9 2.8 2.9
+OfficeHome A. 55.3 53.8 55.5 25.8 28.8 25.3 4.4 5.1 4.7 2.0 5.2 2.5 2.9 3.1 3.5 0.7 1.2 1.0
+C. 49.4 45.5 50.6 27.0 32.6 30.4 15.2 14.9 16.0 6.4 8.9 8.0 12.4 11.2 13.0 6.1 8.3 7.7
+P. 61.7 58.1 62.5 48.0 46.9 51.6 13.2 13.9 14.0 8.6 10.7 10.0 9.2 8.8 10.2 8.3 7.9 8.4
+R. 65.3 63.2 65.6 36.5 40.1 41.0 7.5 7.9 7.9 5.3 9.4 8.9 5.2 4.8 5.6 2.9 2.8 2.9
+DomainNet C. 57.8 50.9 58.8 33.3 34.3 35.0 21.6 18.7 22.8 18.4 19.6 20.0 15.8 12.5 16.6 7.0 7.5 7.8
+I. 35.8 28.0 37.0 12.2 13.3 13.2 6.1 3.7 6.7 4.6 5.3 5.1 3.3 1.9 3.7 0.9 0.9 0.9
+P. 43.9 39.0 44.3 18.4 20.6 20.3 3.1 2.8 3.3 8.6 10.4 9.9 1.8 1.3 1.9 0.3 0.3 0.3
+Q. 12.9 10.3 13.2 10.9 10.8 11.1 8.4 6.8 8.6 5.4 5.4 5.6 7.1 5.4 7.4 4.9 4.8 5.1
+R. 62.1 55.9 62.4 34.5 35.9 36.5 7.1 6.5 7.5 17.6 19.7 19.6 4.5 3.4 4.7 1.2 1.4 1.4
+S. 49.1 43.3 49.7 25.7 26.0 27.5 17.8 15.5 18.6 13.6 14.4 15.1 13.4 10.2 13.9 5.0 5.2 5.6
+Appendix 0.B Data Synthesis
+0.B.1 Synthetic Image Generation
+We employ the stable diffusion [39] v2.1 model for generating synthetic images used in our comparing experiments, specifically utilizing the Stable Diffusion v2-1 Model Card available on Hugging Face [^3] https://huggingface.co/stabilityai/stable-diffusion-2-1.
+For each class across the four datasets, we produce 480 images using our synthetic template prompts as input for the stable diffusion model.
+All generated images are of $512\times 512$ resolution.
+Examples of these synthetic images alongside their corresponding text prompts are displayed in Fig.
 7.
-Figure 7: Representative examples of synthetic images generated using Stable Diffusion v2.1, paired with their associated template-based prompts.
-Table 11: Zero-shot performance on VLCS across varied augmentation combinations and prompt sources.
-We evaluate average top-1 accuracy (%) for each domain using CLIP and CLAP variants trained with different prompt augmentations: ① Object Size Deletion, ② Object Color Deletion, ③ Image Type Deletion, ④ Art Style Deletion, ⑤ Prompt Order Swapping, and ⑥ Gaussian Noise Addition.
-Rightmost columns show EDA-based augmentations: LLM, Random, PromptStyler (Pr.St.), and Template.
-Method Domain Avg. top-1 accuracy (%) (↑) on VLCS
-CLIP (base) ①②③ ①②③ ①②③ ①②③ ③④⑤ ①②⑤ LLM Random Pr.St. Template
-④⑤⑥ ④⑤ ④⑥
-ZS(C) C 99.7 99.9 99.8 99.9 99.8 99.9 99.9 97.9 99.7 99.9 99.9
-L 61.8 66.6 62.3 67.0 62.2 66.2 67.7 66.2 69.0 67.3 66.5
-S 70.1 78.1 75.5 78.0 74.3 78.5 78.0 73.2 76.9 73.5 76.9
-V 73.9 82.8 80.6 83.2 79.3 82.7 84.9 72.6 81.8 81.8 81.9
-ZS(CP) C 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.8 99.9 99.9 99.9
-L 69.9 69.3 67.9 69.6 68.4 70.0 70.4 69.3 70.4 71.2 69.7
-S 73.3 77.6 76.4 76.7 75.9 78.8 77.2 76.2 75.2 75.1 78.0
-V 84.8 85.3 84.0 85.3 84.2 85.1 86.0 77.0 84.2 86.0 84.6
-ZS(PC) C 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9
-L 70.2 70.0 68.0 70.1 68.5 70.0 70.7 67.5 70.6 71.8 70.0
-S 73.6 76.6 75.6 76.0 74.8 77.8 76.9 76.9 75.1 74.9 78.2
-V 86.1 85.7 84.7 85.7 84.5 85.5 86.2 78.2 84.6 86.8 84.8
-ZS(NC) C 87.0 99.8 99.6 99.8 99.4 99.7 99.9 95.3 98.6 99.6 99.8
-L 55.9 65.2 61.3 65.6 60.5 65.4 65.9 63.0 66.7 64.0 64.7
-S 61.4 75.6 70.3 75.2 68.3 74.9 75.3 68.9 73.3 69.8 73.0
-V 68.9 80.1 75.2 80.4 73.8 79.4 82.9 69.3 79.6 77.2 78.6
-Table 12: Ablation study on hyperparameters for CLAP using ViT-B/16 on the VLCS dataset.
-We report domain-wise average top-1 accuracy (%) under three prompt settings—ZS(C), ZS(CP), and ZS(PC)—by varying: (1) $\tau$ (temperature), (2) latent dimension size, and (3) $\alpha$ (residual weight).
-Hyper- Value Average Top-1 Accuracy (%) on VLCS (↑)
+Figure 7: Examples of synthetic images created with SDv2.1 and associated prompts.
+Table 11: Zero-Shot Performance on VLCS Dataset Across Varied Augmentation Combinations and Prompt Sources: ① Random Object Size Deletion, ② Random Object Color Deletion, ③ Random Image Type Deletion, ④ Random Art Style Deletion, ⑤ Random Swapping Order, ⑥ Addition of Gaussian Noise.
+Method Domains Avg. top-1 acc. (%) (↑) of different augmentations and prompts on VLCS
+CLIP (base) ①②③ ①②③ ①②③ ①②③ ③④⑤ ①②⑤ EDA
+④⑤⑥ ④⑤ ④⑥  LLM Rand. Pr.St. Temp.
+ZS(C) C. 99.7 99.9 99.8 99.9 99.8 99.9 99.9 97.9 99.7 99.9 99.9
+L. 61.8 66.6 62.3 67.0 62.2 66.2 67.7 66.2 69.0 67.3 66.5
+S. 70.1 78.1 75.5 78.0 74.3 78.5 78.0 73.2 76.9 73.5 76.9
+V. 73.9 82.8 80.6 83.2 79.3 82.7 84.9 72.6 81.8 81.8 81.9
+ZS(CP) C. 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.8 99.9 99.9 99.9
+L. 69.9 69.3 67.9 69.6 68.4 70.0 70.4 69.3 70.4 71.2 69.7
+S. 73.3 77.6 76.4 76.7 75.9 78.8 77.2 76.2 75.2 75.1 78.0
+V. 84.8 85.3 84.0 85.3 84.2 85.1 86.0 77.0 84.2 86.0 84.6
+ZS(PC) C. 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9 99.9
+L. 70.2 70.0 68.0 70.1 68.5 70.0 70.7 67.5 70.6 71.8 70.0
+S. 73.6 76.6 75.6 76.0 74.8 77.8 76.9 76.9 75.1 74.9 78.2
+V. 86.1 85.7 84.7 85.7 84.5 85.5 86.2 78.2 84.6 86.8 84.8
+ZS(NC) C. 87.0 99.8 99.6 99.8 99.4 99.7 99.9 95.3 98.6 99.6 99.8
+L. 55.9 65.2 61.3 65.6 60.5 65.4 65.9 63.0 66.7 64.0 64.7
+S. 61.4 75.6 70.3 75.2 68.3 74.9 75.3 68.9 73.3 69.8 73.0
+V. 68.9 80.1 75.2 80.4 73.8 79.4 82.9 69.3 79.6 77.2 78.6
+Table 12: Ablative study of hyperparameters on VLCS dataset using ViT-B/16 model.
+Hyper- Value Avg. top-1 acc. (%) (↑) using ViT-B/16 on VLCS dataset
 ZS (C) ZS (CP) ZS (PC)
-C L S V C L S V C L S V
+parameters    C.    L.    S.    V.    C.    L.    S.    V.    C.    L.    S.    V.
 $\tau$ 0.1 99.9 67.6 77.5 84.2 99.9 70.9 74.9 85.9 99.9 71.2 74.6 86.3
@@
 0.9 99.9 66.0 77.6 83.3 99.9 69.0 77.9 85.0 99.9 69.7 77.5 85.0
-Latent Dim. 128 99.9 66.0 77.6 82.6 99.9 70.0 77.4 85.4 99.9 70.1 77.1 85.7
-192 99.9 64.9 77.9 83.0 99.9 68.9 78.0 85.6 99.9 69.0 77.8 86.0
-256 99.9 63.8 77.6 82.7 99.9 67.6 78.7 84.8 99.9 67.8 78.6 85.2
-320 99.9 66.0 77.8 82.9 99.9 69.2 78.1 85.3 99.9 69.7 77.7 85.5
-384 99.9 65.8 76.9 82.8 99.9 69.4 77.5 85.3 99.9 69.6 77.0 85.5
-448 99.9 65.8 77.4 82.1 99.9 69.7 77.6 84.9 99.9 69.9 77.1 85.6
-512 99.9 67.7 78.0 84.9 99.9 70.4 77.2 86.0 99.9 70.7 76.9 86.2
+128.0 99.9 66.0 77.6 82.6 99.9 70.0 77.4 85.4 99.9 70.1 77.1 85.7
+192.0 99.9 64.9 77.9 83.0 99.9 68.9 78.0 85.6 99.9 69.0 77.8 86.0
+256.0 99.9 63.8 77.6 82.7 99.9 67.6 78.7 84.8 99.9 67.8 78.6 85.2
+Lantent 320.0 99.9 66.0 77.8 82.9 99.9 69.2 78.1 85.3 99.9 69.7 77.7 85.5
+dim. 384.0 99.9 65.8 76.9 82.8 99.9 69.4 77.5 85.3 99.9 69.6 77.0 85.5
+448.0 99.9 65.8 77.4 82.1 99.9 69.7 77.6 84.9 99.9 69.9 77.1 85.6
+512.0 99.9 67.7 78.0 84.9 99.9 70.4 77.2 86.0 99.9 70.7 76.9 86.2
 $\alpha$ $10^{-1.5}$ 99.9 66.5 77.9 83.1 99.9 70.4 77.1 86.0 99.9 70.3 76.6 86.1
@@
 $10^{1.5}$ 99.8 72.1 72.8 85.4 99.8 72.2 73.3 85.7 99.8 72.6 72.7 85.5
-B.2 LLM Prompt Generation
+0.B.2 LLM Prompts Generation
 We employ ChatGPT-3.5 [3] to generate the LLM-based prompts used in our comparative analysis of prompt sources.
@@
 For each class, we generate 120 distinct samples.
-table 13 shows representative examples of the generated prompts.
+Tab.
+13 shows representative examples of the generated prompts.
 LLM Prompting Procedure (ChatGPT-3.5) User: I have some class names, such as "dog", "car", "person", "chair".
@@
 A graceful figure skater glides across an ice rink, captured in a delicate, winter-themed pastel drawing.
-B.3 Experiments on ViT-L/14
-We refine the output dimension to match the ViT-L/14 input dimension of 768.
-The latent dimensions are set to 448 for PACS, 640 for VLCS, and 768 for both OfficeHome and DomainNet.
-The inference weighting parameter $\alpha$ is configured as 0.1 for PACS, 0.03 for VLCS, 0.14 for OfficeHome, and 0.2 for DomainNet.
-All other training configurations remain consistent with those used in the ViT-B/16 experiments.
-The training setup for Im.Aug mirrors that of CLAP for each dataset, with $\alpha$ set to 0.1 for PACS and 0.03 for the other three datasets.
-Table 14 presents the zero-shot results for the ViT-L/14 model across four distinct prompts, following the same evaluation protocol as in the ViT-B/16 experiments.
-These results demonstrate that CLAP is more effective than Im.Aug in enhancing zero-shot performance.
-Furthermore, table 15 shows that CLAP substantially reduces performance variance across prompts, further validating its improvements over CLIP across multiple model scales.
-Detailed domain-level results are reported in table 16, providing a comprehensive breakdown of performance by domain.
-B.4 Experiments on ResNet50x16
-To validate our approach on different model structures, we repeated zero-shot experiments on the ResNet50x16 model pre-trained with CLIP.
-Since the output dimension of CLIP is the same as ViT-B/16, we used the same training configuration as ViT-B/16 for training Im.Aug and CLAP.
-For inference, we refined the weighting coefficient $\alpha$ to 0.1, 1, 0.03, and 0.1 for Im.Aug, and 0.03, 0.2, 0.06, and 0.1 for CLAP, for PACS, VLCS, OfficeHome, and DomainNet respectively.
-Table 17 showcases the zero-shot results for ResNet50x16 model across different prompts, substantiating that CLAP is more effective than Im.Aug in refining CLIP features.
-Moreover, table 18 illustrates that both Im.Aug and CLAP reduce variations in zero-shot performance across different prompts, with the improvement of CLAP being more significant.
-The results validate our approach across different model scales, including both ViT-based and CNN-based structures.
-Domain-level results are detailed in table 19.
-Appendix C Discussion
-C.1 Rationale behind CLAP’s Foundation on CLIP
-The primary challenge in cross-modal transferability lies in the significant domain gap between text and image data, which typically hinders the direct application of models trained in one modality to another.
-For a causal explaination, despite the consistency of the content variable that dictates the object label across modalities, the generative processes from latent variables to observations inherent to each modality differ markedly.
-The CLIP model, trained on a comprehensive dataset of image-text pairs with a symmetric InfoNCE loss, significantly ameliorates this issue.
-By aligning the features of text and images into similar patterns, it facilitates leveraging a network trained atop the CLIP encoder of one modality as a viable proxy for the other.
-Consequently, this allows for the direct application of the disentangled network trained in the text modality atop CLIP’s image encoder to refine representations.
-C.2 Impact of Image and Text Augmentations
-Identifying pure content factors poses a significant challenge.
-This difficulty primarily arises from the need for finding effective augmentations of observational data to alter style factors significantly while preserving content integrity.
-Through the cross-modal alignment provided by CLIP, we discovered that disentangling in one modality can seamlessly improve representations in both modalities.
-The impact of image augmentations has been well-explored and found effective at preserving content, but traditional methods do not impose sufficient changes to remove all style information.
-Our exploration of text augmentations reveals that the logical structure of text and the relative ease of implementing style changes can have a significant impact on achieving disentanglement.
-However, more efficient methods are worthy of exploration.
-A promising direction for future research is to explore efficient combinations of both modalities to enhance disentangled semantics.
-As each modality has its unique advantages—Text data recapitulates properties well since it is pre-processed by human intelligence, while image data is more precise in depicting the exact same objects or events due to its more detailed nature—the impact of combining augmentations of both modalities could be substantial.
+Appendix 0.C Experiments on Other CLIP Model Scales
+0.C.1 Experiments on ViT-L/14
+We refined the output dimension to align with the input dimension of 768.
+The chosen latent dimensions were 448 and 640 for PACS and VLCS, respectively, and 768 for both OfficeHome and DomainNet.
+The inference weighting $\alpha$ was set to 0.1 for PACS, 0.03 for VLCS, 0.14 for OfficeHome, and 0.2 for DomainNet.
+All other training configurations remained consistent with the ViT-B/16 experiments across each dataset.
+The training configuration for Im.Aug was set the same as CLAP for each dataset, with the inference weighting $\alpha$ being 0.1 for PACS and 0.03 for the other three datasets.
+Table 14 showcases the zero-shot results for the ViT-L/14 model using four distinct prompts, following the protocol established for the ViT-B/16 experiments.
+These results demonstrate that CLAP is more efficient than Im.Aug in enhancing zero-shot performance.
+Moreover, Tab.
+15 illustrates that CLAP significantly reduces variations in zero-shot performance across different prompts, thereby confirming CLAP’s performance improvements over CLIP across a range of model sizes.
+Detailed domain-level results are presented in Tab.
+16, offering an in-depth analysis.
 Table 14: Zero-shot performance across four prompts ("C", "PC", "CP") and 1 noised prompts ("NC") with CLIP pre-trained ViT-L/14 model.
@@
 CLAP 98.5 73.1 81.3 58.3 77.8
-Table 15: Prompt robustness analysis with CLIP pre-trained ViT-L/14 model. CLAP reduces the variance in zero-shot performance across different prompts.
+Table 15: CLAP reduces the variance in zero-shot performance across different prompts with CLIP pre-trained ViT-L/14 model.
 Metric Method Zero-shot variance, avg. top-1 acc. (%) (↓)
@@
 CLAP 0.1 7.7 6.3 5.9 5.0
-Table 16: Domain-level zero-shot results for the ViT-L/14 model across test datasets.
-CLAP consistently improves performance across all domains and prompt types, demonstrating its robustness and generalization capabilities.
+Table 16: Domain-level zero-shot results of the ViT-L/14 model on the test datasets.
 Datasets Domains Domain-level avg. top-1 acc. (%) of zero-shot performance using ViT-L/14 (↑)
@@
 V. 80.0 81.6 84.3 85.2 85.7 86.2 85.1 85.3 86.4 65.0 69.3 74.3
-O.H. A. 86.2 86.3 87.7 85.7 86.2 88.1 87.0 87.0 87.8 78.1 77.1 80.7
+OfficeHome A. 86.2 86.3 87.7 85.7 86.2 88.1 87.0 87.0 87.8 78.1 77.1 80.7
 C. 73.3 73.4 75.7 73.8 73.4 76.0 73.1 73.5 76.0 65.9 66.3 70.6
@@
 R. 92.2 92.7 93.0 92.2 92.4 93.4 93.1 93.3 93.9 83.8 84.0 86.9
-D.N. C. 78.4 78.5 79.1 77.5 77.7 78.8 79.4 79.4 79.7 70.0 70.4 72.8
+DomainNet C. 78.4 78.5 79.1 77.5 77.7 78.8 79.4 79.4 79.7 70.0 70.4 72.8
 I. 52.9 53.0 54.6 50.4 50.7 53.6 51.7 52.0 53.9 45.3 45.2 48.8
@@
 S. 70.2 70.4 70.7 69.4 69.8 70.6 71.0 71.3 71.5 62.0 62.2 64.6
-Table 17: Zero-shot performance with the CLIP pre-trained ResNet50x16 model. CLAP consistently enhances performance across all datasets, validating its effectiveness.
+0.C.2 Experiments on ResNet50x16
+To validate our approach on different model structures, we repeated zero-shot experiments on the ResNet50x16 model pre-trained with CLIP.
+Since the output dimension of CLIP is the same as ViT-B/16, we used the same training configuration as ViT-B/16 for training Im.Aug and CLAP.
+For inference, we refined the weighting coefficient $\alpha$ to 0.1, 1, 0.03, and 0.1 for Im.Aug, and 0.03, 0.2, 0.06, and 0.1 for CLAP, for PACS, VLCS, OfficeHome, and DomainNet respectively.
+Table 17 showcases the zero-shot results for ResNet50x16 model across different prompts, substantiating that CLAP is more effective than Im.Aug in refining CLIP features.
+Moreover, Tab.
+18 illustrates that both Im.Aug and CLAP reduce variations in zero-shot performance across different prompts, with the improvement of CLAP being more significant.
+The results validate our approach across different model scales, including both ViT-based and CNN-based structures.
+Domain-level results are detailed in Tab.
+19.
+Table 17: Zero-shot performance with CLIP pre-trained ResNet50x16 model. CLAP demonstrates consistent enhancement across all datasets, validating its effectiveness.
 Prompt Method Zero-shot performance, avg. top-1 acc. (%) (↑)
@@
 CLAP 94.9 80.1 71.9 50.6 74.4
-Table 18: CLAP consistently reduces variance in zero-shot performance across different prompts using the CLIP pre-trained ResNet50x16 model, further validating its robustness and effectiveness.
+Table 18: CLAP consistently reduces variances in zero-shot performance across different prompts with CLIP pre-trained ResNet50x16 model, validating its effectiveness.
 Metric Method Zero-shot variance, avg. top-1 acc. (%) (↓)
@@
 CLAP 2.1 -0.1 9.7 7.5 4.8
-Table 19: Domain-level zero-shot performance using ResNet50x16 across test datasets.
-CLAP consistently improves upon CLIP and Im.Aug across domains and prompt types (ZS(C), ZS(CP), ZS(PC), ZS(NC)), demonstrating its generalizability and robustness.
+Table 19: Domain-level zero-shot results using RestNet50x16 on the test datasets.
 Datasets Domains Domain-level avg. top-1 acc. (%) of zero-shot performance using RN50x16 (↑)
@@
 V. 68.4 70.1 85.2 74.5 72.1 85.3 81.7 78.0 85.2 63.1 68.5 84.5
-O.H. A. 82.2 82.5 83.5 79.7 79.9 80.6 82.0 82.4 83.4 67.7 68.9 72.1
+OfficeHome A. 82.2 82.5 83.5 79.7 79.9 80.6 82.0 82.4 83.4 67.7 68.9 72.1
 C. 63.0 62.9 64.7 61.7 62.2 62.8 65.4 65.3 66.1 54.6 55.0 56.8
@@
 R. 88.1 88.2 89.1 87.3 87.5 87.6 89.2 89.5 89.7 79.5 78.9 80.3
-D.N. C. 69.0 68.9 69.6 68.6 68.6 69.4 69.9 70.0 70.4 59.5 60.1 61.4
+DomainNet C. 69.0 68.9 69.6 68.6 68.6 69.4 69.9 70.0 70.4 59.5 60.1 61.4
 I. 51.0 51.1 52.7 48.2 49.0 50.6 48.2 48.9 50.7 41.2 41.6 44.3
@@
 S. 63.2 63.1 63.6 62.0 62.4 63.4 63.9 63.8 64.6 53.1 53.4 55.3
+Appendix 0.D Discussion
+0.D.1 Rationale behind CLAP’s Foundation on CLIP
+The primary challenge in cross-modal transferability lies in the significant domain gap between text and image data, which typically hinders the direct application of models trained in one modality to another.
+For a causal explaination, despite the consistency of the content variable that dictates the object label across modalities, the generative processes from latent variables to observations inherent to each modality differ markedly.
+The CLIP model, trained on a comprehensive dataset of image-text pairs with a symmetric InfoNCE loss, significantly ameliorates this issue.
+By aligning the features of text and images into similar patterns, it facilitates leveraging a network trained atop the CLIP encoder of one modality as a viable proxy for the other.
+Consequently, this allows for the direct application of the disentangled network trained in the text modality atop CLIP’s image encoder to refine representations.
+0.D.2 Impact of Image and Text Augmentations
+Identifying pure content factors poses a significant challenge.
+This difficulty primarily arises from the need for finding effective augmentations of observational data to alter style factors significantly while preserving content integrity.
+Through the cross-modal alignment provided by CLIP, we discovered that disentangling in one modality can seamlessly improve representations in both modalities.
+The impact of image augmentations has been well-explored and found effective at preserving content, but traditional methods do not impose sufficient changes to remove all style information.
+Our exploration of text augmentations reveals that the logical structure of text and the relative ease of implementing style changes can have a significant impact on achieving disentanglement.
+However, more efficient methods are worthy of exploration.
+A promising direction for future research is to explore efficient combinations of both modalities to enhance disentangled semantics.
+As each modality has its unique advantages—Text data recapitulates properties well since it is pre-processed by human intelligence, while image data is more precise in depicting the exact same objects or events due to its more detailed nature—the impact of combining augmentations of both modalities could be substantial.